### v1

# Analytic and Variational Stability of Deep Learning Systems

Abstract
We propose a unified analytic and variational framework for studying stability in deep learning systems viewed as coupled representation–parameter dynamics. The central object is the Learning Stability Profile, which tracks the infinitesimal response of representations, parameters, and update mechanisms to perturbations along the learning trajectory. We prove a Fundamental Analytic Stability Theorem showing that uniform boundedness of these stability signatures is equivalent, up to norm equivalence, to the existence of a Lyapunov-type energy that dissipates along the learning flow.
In smooth regimes, the framework yields explicit stability exponents linking spectral norms, activation regularity, step sizes, and learning rates to contractivity of the learning dynamics. Classical spectral stability results for feedforward networks, a discrete CFL-type condition for residual architectures, and parametric and temporal stability laws for stochastic gradient methods arise as direct consequences. The theory extends to non-smooth learning systems, including ReLU networks, proximal and projected updates, and stochastic subgradient flows, by replacing classical derivatives with Clarke generalized derivatives and smooth energies with variational Lyapunov functionals.
The resulting framework provides a unified dynamical description of stability across architectures and optimization methods, clarifying how architectural and algorithmic choices jointly govern robustness and sensitivity to perturbations. It also provides a foundation for further extensions to continuous-time limits and geometric formulations of learning dynamics.
Keywords Analytic stability $\cdot$ learning dynamics $\cdot$ Lyapunov methods $\cdot$ energy-dissipative systems $\cdot$ generalization theory $\cdot$ deep neural networks $\cdot$ dynamical systems.
1 Introduction
Stability plays a central role in learning theory, but is formalized differently across fields. In numerical analysis, stability refers to bounded responses of discrete evolutions to perturbations and is certified by Lyapunov energies [HirschSmaleDevaney2012, Strikwerda2004]. In statistical learning theory, stability is defined through algorithmic sensitivity to data perturbations, beginning with [Bousquet2002] and extended to stochastic gradient methods by [Hardt2016Train]. A separate literature studies stability through Jacobian spectra and curvature using tools from dynamical systems and control [Sontag1998]. Recent empirical work further shows that explanations and attribution scores of modern networks can be highly unstable under perturbations of inputs, parameters, or training conditions [Fel2022, Gan2025].
Despite their maturity, these perspectives rely on distinct primitives and do not provide a unified mathematical description of stability in deep learning. This paper introduces a single analytic and variational framework that unifies these views.
The central construct is the Learning Stability Profile (LSP), which records how infinitesimal perturbations propagate through the coupled representation, parameter, and update dynamics of learning. The LSP is defined at the level of the full learning flow and applies uniformly to smooth and non-smooth architectures, including ReLU networks, proximal and projected methods, stochastic subgradient flows, residual networks, and continuous-time limits. Our main result shows that, up to equivalence of norms, boundedness of the Learning Stability Profile is equivalent to the existence of a Lyapunov-type energy that dissipates along the learning trajectory.
This equivalence yields concrete stability laws. In smooth regimes it recovers conditions based on Jacobian bounds, spectral regularity, and step-size constraints. For residual networks it produces a discrete Courant–Friedrichs–Lewy type restriction consistent with explicit time stepping [Strikwerda2004]. In stochastic and non-smooth regimes, the same analytic structure yields stability statements consistent with stochastic approximation theory [KushnerYin2003, Benaim1999] and generalized variational analysis [Clarke1990Optimization, RockafellarWets1998].
The contribution is therefore not an architecture-specific condition, but a unified analytic principle connecting perturbation growth, stability exponents, and variational energy dissipation across modern learning systems.
2 Setup
We model learning as a coupled discrete-time dynamical system with state
$Z_{k,t}=(X_{k,t},\theta_{k,t},U_{k,t}),$
where $X_{k,t}$ denotes representations, $\theta_{k,t}$ parameters, and $U_{k,t}$ update variables such as optimization states or stochastic noise.
The system evolves through three maps: a representation map $F_{\theta}$, a parameter update map $\mathcal{A}$, and an update dynamics map $\mathcal{U}$. A single learning step is
$X_{k+1,t}=F_{\theta_{k,t}}(X_{k,t}),\qquad\theta_{k,t+1}=\mathcal{A}(\theta_{k,t},U_{k,t}),\qquad U_{k,t+1}=\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t}).$
Derivatives are classical when the maps are smooth and understood in the Clarke sense when they are only locally Lipschitz [Clarke1990Optimization]. When updates involve randomness, the system is interpreted as a measurable stochastic process [KushnerYin2003, Benaim1999].
To quantify sensitivity, we consider perturbations of the initial state $Z_{0,0}$. The first-order response of a single learning step admits the block representation
$\mathcal{J}_{k,t}=\begin{pmatrix}D_{X}F_{\theta}&D_{\theta}F_{\theta}&0\\ 0&D_{\theta}\mathcal{A}&D_{U}\mathcal{A}\\ D_{X}\mathcal{U}&D_{\theta}\mathcal{U}&D_{U}\mathcal{U}\end{pmatrix}$,
interpreted as a classical or generalized Jacobian. The sequence $(\mathcal{J}_{k,t})$ constitutes the Learning Stability Profile and encodes perturbation propagation jointly across depth and time.
Asymptotic growth of these sensitivities is summarized by analytic stability exponents. For representations,
$\alpha_{x}=\limsup_{n\to\infty}\frac{1}{n}\log\bigl\|D_{Z_{0,0}}X_{n,0}\bigr\|,$
with analogous definitions for $\alpha_{\theta}$ and $\alpha_{u}$. Nonpositive exponents correspond to stability and strictly negative exponents to strict stability [HirschSmaleDevaney2012].
A variational learning energy is a function $\mathcal{E}(Z)$ satisfying a dissipation inequality
$\mathcal{E}(Z_{k+1,t})-\mathcal{E}(Z_{k,t})\leq-\gamma\,\Psi(Z_{k,t}),$
for some $\gamma>0$ and nonnegative $\Psi$. This generalizes discrete Lyapunov theory in the variational framework of [RockafellarWets1998]. The next section formalizes the Learning Stability Profile and establishes the analytic stability theorem.
Definition 2.1 (Learning System).
A learning system is specified by maps
$F_{\theta}:\mathcal{X}_{k}\to\mathcal{X}_{k+1},\qquad\mathcal{A}:\Theta\times\mathcal{U}\to\Theta,\qquad\mathcal{U}:\mathcal{U}\times\mathcal{X}_{0}\times\Theta\to\mathcal{U},$
which induce the joint state evolution
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t})=\bigl(F_{\theta_{k,t}}(X_{k,t}),\,\mathcal{A}(\theta_{k,t},U_{k,t}),\,\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t})\bigr)$
on the product space
$\mathcal{M}=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta\times\mathcal{U}.$
A trajectory is any sequence $(Z_{k,t})$ obtained by iterating $\Phi_{k,t}$ from admissible initial conditions.
Assumption 1 (Analytic Regularity).
Each map $\Phi_{k,t}$ is either continuously differentiable or locally Lipschitz with a nonempty Clarke generalized Jacobian. Constants
$L_{\mathrm{fwd}},\qquad L_{\mathrm{par}},\qquad L_{\mathrm{upd}}$
exist such that for all $(x,\theta,u)$, and every Jacobian or generalized Jacobian element
$V\in\partial\Phi_{k,t}(x,\theta,u),$
the corresponding blocks satisfy
$\|V_{x}\|\leq L_{\mathrm{fwd}},\qquad\|V_{\theta}\|\leq L_{\mathrm{par}},\qquad\|V_{u}\|\leq L_{\mathrm{upd}}.$
This ensures uniformly bounded directional derivatives in the representation, parameter, and update directions.
3 Learning Stability Profile
Section 2 models learning as a discrete-time dynamical system on a joint state space. At layer index $k$ and training step $t$, the state is
$Z_{k,t}=(X_{k,t},\theta_{k,t},U_{k,t}),$
and one step of the learning flow is
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t})=\bigl(F_{\theta_{k,t}}(X_{k,t}),\,\mathcal{A}(\theta_{k,t},U_{k,t}),\,\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t})\bigr).$
By Assumption 1, each $\Phi_{k,t}$ is either continuously differentiable or locally Lipschitz with generalized derivatives in the sense of [Clarke1990Optimization]. This section defines the Learning Stability Profile, which quantifies the propagation of infinitesimal perturbations in the input, parameter, and update directions.
3.1 Directional sensitivities of the learning flow
Define the product manifold
$\mathcal{M}=\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\times\mathcal{U},$
equipped with any compatible norm. Since $\mathcal{M}$ is finite dimensional, all such norms are equivalent.
Fix an architecture and training protocol. Consider three perturbations: an input perturbation $\delta x$ of $X_{0,0}$, an initialization perturbation $\delta\theta^{0}$ of $\theta_{0,0}$, and an update perturbation $\delta u$ of the update sequence $(U_{k,t})$. Under Assumption 1, the dependence of $Z_{k,t}$ on these perturbations is differentiable in the classical or Clarke sense [Clarke1990Optimization].
The associated directional sensitivity operators are
$\mathsf{S}^{x}_{k,t}=\frac{\partial Z_{k,t}}{\partial x}\Big|_{(x_{0},\theta^{0},u)},\qquad\mathsf{S}^{\theta}_{k,t}=\frac{\partial Z_{k,t}}{\partial\theta^{0}}\Big|_{(x_{0},\theta^{0},u)},\qquad\mathsf{S}^{u}_{k,t}=\frac{\partial Z_{k,t}}{\partial u}\Big|_{(x_{0},\theta^{0},u)}.$
In the smooth case these coincide with classical derivatives of the composite map $(x,\theta^{0},u)\mapsto Z_{k,t}$. In the non-smooth case they are elements of the Clarke generalized Jacobian [Clarke1990Optimization, RockafellarWets1998].
The collection
$\bigl\{\mathsf{S}^{x}_{k,t},\,\mathsf{S}^{\theta}_{k,t},\,\mathsf{S}^{u}_{k,t}\bigr\}_{k,t}$
constitutes the Learning Stability Profile and encodes linearized perturbation propagation across depth and training time.
3.2 Analytic stability exponents
We now compress the local stability signatures into asymptotic exponents. These quantities measure the effective growth or decay of infinitesimal perturbations along both depth and training time, and play a role analogous to Lyapunov exponents in dynamical systems [HirschSmaleDevaney2012].
Definition 3.1 (Analytic stability exponents).
For a fixed architecture and training protocol, let $L$ and $T$ denote network depth and training horizon. The analytic forward exponent is
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\Bigr).$
The analytic parameter exponent is
$\alpha_{\theta}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{\theta}_{k,t}\Bigr),$
and the analytic update exponent is
$\alpha_{u}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{u}_{k,t}\Bigr).$
These exponents are well defined under Assumption 1, since the learning flow is locally Lipschitz and the stability signatures are finite along any finite horizon. They represent large-scale growth rates of infinitesimal perturbations. Negative values correspond to exponential decay, nonpositive values to uniform boundedness, and positive values to instability.
Analytic stability is expressed as constraints on these exponents.
Definition 3.2 (Analytic stability and stability indices).
The learning system is analytically stable if
$\alpha_{x}\leq 0,\qquad\alpha_{\theta}\leq 0,\qquad\alpha_{u}\leq 0.$
It is strictly analytically stable if all three exponents are strictly negative. A triple $(\beta_{x},\beta_{\theta},\beta_{u})$ with $\beta_{x},\beta_{\theta},\beta_{u}\leq 0$ is called a stability index if
$\alpha_{x}\leq\beta_{x},\qquad\alpha_{\theta}\leq\beta_{\theta},\qquad\alpha_{u}\leq\beta_{u}.$
Definition 3.3 (Analytic learning energy).
A functional
$\mathcal{E}:\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}$
is called an analytic learning energy if it satisfies the following conditions.
  1. 1.
    Coercivity. There exist constants $c_{1},c_{2}>0$ such that, for all admissible states,
    $c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}(X_{0},\dots,X_{L},\theta)\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr).$
  2. 2.
    Differentiability. The map $\mathcal{E}$ is continuously Fréchet differentiable with respect to each of its arguments.
  3. 3.
    Compatibility with the learning flow. For every layer $k$, the directional derivatives $\nabla_{X_{k}}\mathcal{E}$ and $\nabla_{\theta}\mathcal{E}$ exist and are well defined along the learning trajectory generated by the update maps $f_{k}$ and $\mathcal{A}_{t}$.
A functional satisfying these properties couples the discrete learning dynamics to a smooth Lyapunov structure.
Assumption 2 (Analytic energy–dissipation law).
There exist an analytic learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$, linear at first order in the initial perturbations $(\delta x,\delta\theta^{0},\delta u)$, such that for every training step $t$ and every admissible trajectory,
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R\bigl(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u\bigr).$
The constant $\gamma$ is the dissipation rate. The remainder satisfies the uniform bound
$|R|\leq C\bigl(\|\delta x\|+\|\delta\theta^{0}\|+\|\delta u\|\bigr)$
for some $C>0$ independent of depth, training horizon, and state.
The following basic properties make these definitions robust.
Proposition 3.4 (Independence of norm).
Let $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ be two norms on $\mathcal{M}$ inducing operator norms $\|\cdot\|_{1,\mathrm{op}}$ and $\|\cdot\|_{2,\mathrm{op}}$, associated stability signatures $\sigma^{x,1}_{k,t}$, $\sigma^{x,2}_{k,t}$, and exponents $\alpha_{x}^{(1)}$, $\alpha_{x}^{(2)}$. Then there exist constants $c,C>0$ such that for all $k,t$
$c\,\sigma^{x,1}_{k,t}\leq\sigma^{x,2}_{k,t}\leq C\,\sigma^{x,1}_{k,t},$
and the exponents coincide, $\alpha_{x}^{(1)}=\alpha_{x}^{(2)}$. The same conclusion holds for $\alpha_{\theta}$ and $\alpha_{u}$.
Proof.
Since $\mathcal{M}$ is finite dimensional, the norms $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ are equivalent, so there exist $c,C>0$ such that
$c\|v\|_{1}\leq\|v\|_{2}\leq C\|v\|_{1}\qquad\text{for all }v\in\mathcal{M}.$
This implies corresponding bounds on the induced operator norms and hence on the stability signatures. Taking logarithms, dividing by $L+T$, and passing to the limsup removes additive constants, yielding equality of the exponents. ∎
Proposition 3.5 (Finite horizon bound and exponent sign).
The following statements hold.
  1. 1.
    If there exists a constant $C_{x}\geq 1$ such that $\sigma^{x}_{k,t}\leq C_{x}$ for all $k,t$, then $\alpha_{x}\leq 0$. Conversely, if $\alpha_{x}<0$, then there exist $\varepsilon>0$ and $C^{\prime}_{x}\geq 1$ such that for all sufficiently large $L+T$,
    $\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\leq C^{\prime}_{x}\,\mathrm{e}^{-\varepsilon(L+T)}.$
  2. 2.
    The same statements hold with $\sigma^{\theta}_{k,t}$ and $\alpha_{\theta}$, and with $\sigma^{u}_{k,t}$ and $\alpha_{u}$.
Proof.
If $\sigma^{x}_{k,t}\leq C_{x}$ uniformly, then for all $L,T$
$\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)\leq\frac{1}{L+T}\log C_{x},$
and taking the limsup gives $\alpha_{x}\leq 0$. Conversely, if $\alpha_{x}<0$, there exist $\varepsilon>0$ and $N$ such that for all $L+T\geq N$,
$\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)\leq-\varepsilon.$
Rearranging yields the stated exponential bound, with $C^{\prime}_{x}$ chosen to absorb the finitely many cases $L+T<N$. The arguments for $\alpha_{\theta}$ and $\alpha_{u}$ are identical. ∎
These results show that analytic stability may be characterized either by uniform boundedness of sensitivity operators along the learning trajectory or by nonpositivity of the associated asymptotic exponents. Later sections establish that these equivalent conditions are, in turn, equivalent to the existence of Lyapunov-type energies satisfying discrete dissipation inequalities. This links the local, infinitesimal profile developed here to the global energy structures of classical stability theory [HirschSmaleDevaney2012, RockafellarWets1998].
4 Analytic Stability Theorem
We now connect the Learning Stability Profile of Section 3 with Lyapunov-type energy functionals. Throughout this section we work under Assumption 1 and use the notation
$\mathcal{M}:=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta,$
with norms on product spaces understood to be equivalent Euclidean norms.
The stability exponents of Definition 3.1 summarize the worst-case growth of the three components of the Learning Stability Profile. The following theorem shows that nonpositive exponents are equivalent to the existence of a dissipative analytic learning energy in the sense of Definition 3.3 and Assumption 2. The equivalence holds up to a change of norm on $\mathcal{M}$ and requires no convexity or stationarity beyond Assumption 1.
Theorem 4.1 (Fundamental Analytic Stability Theorem).
Consider a learning system in the sense of Definition 2.1. Suppose that Assumption 1 holds and that the Learning Stability Profile of Definition 3.1 is well defined. Then, up to equivalence of norms on $\mathcal{M}$, the following statements are equivalent.
  1. 1.
    There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all depths $0\leq k\leq L$, all times $0\leq t\leq T$, and all horizons $L,T$,
    $\sigma^{x}_{k,t}\leq C_{x},\qquad\sigma^{\theta}_{k,t}\leq C_{\theta},\qquad\sigma^{u}_{k,t}\leq C_{u}.$
    Equivalently,
    $\alpha_{x}\leq 0,\qquad\alpha_{\theta}\leq 0,\qquad\alpha_{u}\leq 0.$
  2. 2.
    There exist an analytic learning energy $\mathcal{E}$ in the sense of Definition 3.3 and a constant $\gamma>0$ such that every trajectory of the coupled learning flow satisfies the dissipation inequality of Assumption 2,
    $\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$
    where $R$ is first order in $(\delta x,\delta\theta^{0},\delta u)$.
Moreover, when these conditions hold there exist constants $c_{1},c_{2}>0$, depending only on the regularity bounds in Assumption 1, such that the stability indices of Definition 3.2 may be chosen with
$\beta_{x}\leq-c_{1}\gamma,\qquad\beta_{\theta}\leq-c_{1}\gamma,\qquad\beta_{u}\leq-c_{2}\gamma.$
Sketch of proof.
We work up to equivalence of norms on the finite dimensional space $\mathcal{M}$.
Assume first that item (1) holds. The uniform derivative bounds imply that each one-step map $S^{t}\mapsto S^{t+1}$ is globally Lipschitz under a suitable product norm, with constant controlled by $(C_{x},C_{\theta},C_{u})$. The system is therefore incrementally stable in the sense of [Angeli2002Incremental] and admits a smooth Lyapunov function that decays along differences of trajectories. Standard converse Lyapunov results for discrete-time systems yield a coercive $C^{1}$ functional $\mathcal{E}$ on $\mathcal{M}$ satisfying
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\|S^{t}_{\cdot}\|^{2}+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$
for some $\gamma>0$ and a perturbation remainder linear at first order. This $\mathcal{E}$ is an analytic learning energy, establishing item (2).
Assume now that item (2) holds. Linearizing the learning flow and the energy $\mathcal{E}$ along an arbitrary trajectory yields a discrete-time Lyapunov inequality for the linearized system. If the differential response in any of the $(x,\theta,u)$ directions were unbounded, this inequality would be violated for sufficiently large time. A contradiction argument using standard discrete-time Lyapunov theory [Khalil2002Nonlinear, Chapter 4] yields uniform bounds on the operator norms defining $\sigma^{x}_{k,t}$, $\sigma^{\theta}_{k,t}$, and $\sigma^{u}_{k,t}$. Hence the exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ are nonpositive and item (1) follows. The bounds on $(\beta_{x},\beta_{\theta},\beta_{u})$ follow from norm equivalence and exponential decay induced by $\gamma$. ∎
Theorem 4.1 admits three natural projections, each isolating one coordinate of the Learning Stability Profile and its associated exponent.
Corollary 4.2 (Forward analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The forward exponent satisfies $\alpha_{x}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{x}\geq 1$ such that for every input perturbation $\delta x$, every depth $k$ and every time $t$,
    $\|\delta X^{t}_{k}\|\leq C_{x}\|\delta x\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta x$ at first order.
Corollary 4.3 (Parametric analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The parametric exponent satisfies $\alpha_{\theta}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{\theta}\geq 1$ such that for every initialization perturbation $\delta\theta^{0}$, every depth $k$ and every time $t$,
    $\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{\theta}\|\delta\theta^{0}\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta\theta^{0}$ at first order.
Corollary 4.4 (Temporal analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The temporal exponent satisfies $\alpha_{u}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{u}\geq 1$ such that for every admissible perturbation of the update mechanism $\delta u$, and for all depths $k$ and times $t$,
    $\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{u}\|\delta u\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta u$ at first order.
These corollaries show that forward robustness, parametric robustness, and temporal robustness are not independent notions. Each is a coordinate projection of the same analytic object, namely the Learning Stability Profile, and each is equivalent to the existence of a Lyapunov-type energy that dissipates along the corresponding perturbation direction.
5 Residual and Feedforward Stability Laws
We now specialise the analytic stability theory to standard architectures. We consider feedforward networks with spectral norm control and residual networks with an explicit step size. The purpose is not to model all implementation details, but to extract simple analytic laws that recover and sharpen common stability heuristics.
5.1 Feedforward networks with spectral control
Consider an $L$-layer feedforward network of the form
$X_{k+1}=f_{k}(X_{k})=\sigma(W_{k}X_{k}+b_{k}),\qquad k=0,\dots,L-1,$
where $W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}}$ and $b_{k}\in\mathbb{R}^{d_{k+1}}$. Assume that $\sigma$ is applied elementwise, is $1$-Lipschitz, and is differentiable almost everywhere. ReLU, leaky ReLU with slope in $[0,1]$, and hard $\tanh$ satisfy these properties.
At differentiable points the Jacobian has the form
$Jf_{k}(x)=D_{k}(x)W_{k},$
where $D_{k}(x)$ is diagonal with entries in $[0,1]$. Consequently,
$\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}\quad\text{for all differentiable points }x.$
Let $F_{\theta}:=f_{L-1}\circ\cdots\circ f_{0}$, and let $\mathsf{S}^{x}_{L}$ denote the forward sensitivity operator in the sense of Definition 3.1. By the chain rule,
$JF_{\theta}(x)=Jf_{L-1}(X_{L-1})\cdots Jf_{0}(X_{0}),$
and therefore
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|W_{k}\|_{2}.$
Theorem 5.1 (Forward spectral stability for feedforward networks).
Consider the feedforward architecture above and suppose there exists $\rho\in(0,1)$ such that
$\|W_{k}\|_{2}\leq\rho\quad\text{for all }k=0,\dots,L-1.$
Then
$\sup_{x\in\mathcal{X}_{0}}\ \|JF_{\theta}(x)\|_{2}\leq\rho^{L},$
and the forward analytic exponent satisfies
$\alpha_{x}\leq\log\rho<0.$
In particular, the forward component of the Learning Stability Profile is uniformly bounded, and the hypotheses of Theorem 4.1 hold with a quadratic energy depending only on $X_{L}$ and $\theta$.
Sketch of proof.
The Jacobian bound implies that the forward Lipschitz constant of each layer does not exceed $\rho$. Applying the chain rule yields
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|Jf_{k}(X_{k})\|_{2}\leq\rho^{L}.$
This gives the uniform bound on the forward sensitivity and implies that the forward stability signatures decay exponentially with depth. The exponent bound follows directly from the definition of $\alpha_{x}$, and existence of a compatible energy follows from Theorem 4.1. ∎
Theorem 5.1 formalises the standard intuition that spectral norm control ensures stability of deep feedforward networks [Bartlett2017Spectrally, Yoshida2017]. Within the analytic stability framework, the forward exponent $\alpha_{x}$ acts as a depth-normalised logarithmic Lipschitz constant, and the condition $\rho<1$ enforces strict contraction along the representation direction.
5.2 Residual networks and CFL type stability
Residual architectures introduce an explicit step size that links depthwise propagation with time discretisations of continuous dynamical systems [He2016ResNet, HaberRuthotto2017]. Within the analytic stability framework this connection becomes quantitative through the forward stability exponent.
Consider a residual layer
$X_{k+1}=X_{k}+h\,g_{k}(X_{k};\theta_{k}),$ (5.1)
where $h>0$ is a fixed step size and $g_{k}$ is differentiable in $x$ with Jacobian
$G_{k}(x;\theta_{k}):=\frac{\partial g_{k}}{\partial x}(x;\theta_{k}).$
Assume that $g_{k}$ is globally Lipschitz in $x$ with constant $M_{g}$ and uniformly dissipative with rate $m>0$, meaning
$\frac{G_{k}(x;\theta_{k})+G_{k}(x;\theta_{k})^{\top}}{2}\preceq-mI,\qquad\|G_{k}(x;\theta_{k})\|_{2}\leq M_{g}$ (5.2)
for all admissible $(x,\theta_{k})$ and all $k$. Such conditions are standard in the analysis of dissipative ODEs and their explicit time discretisations [Strikwerda2004, HaberRuthotto2017].
At a fixed layer the Jacobian with respect to $X_{k}$ is
$\frac{\partial X_{k+1}}{\partial X_{k}}=I+h\,G_{k}(X_{k};\theta_{k}).$
Its spectral norm can be bounded explicitly in terms of $(m,M_{g})$.
Lemma 5.2 (One step residual Lipschitz factor).
Under assumption (5.2), every residual layer satisfies
$\Bigl\|\frac{\partial X_{k+1}}{\partial X_{k}}\Bigr\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2}$
for all admissible $(X_{k},\theta_{k})$.
Proof.
Let $v$ be a unit vector. Then
$\|(I+hG_{k})v\|_{2}^{2}=\|v\|_{2}^{2}+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
The dissipativity assumption yields $v^{\top}G_{k}v\leq-m\|v\|_{2}^{2}$, while the Lipschitz bound gives $\|G_{k}v\|_{2}^{2}\leq M_{g}^{2}\|v\|_{2}^{2}$. Substitution gives
$\|(I+hG_{k})v\|_{2}^{2}\leq(1-2hm+h^{2}M_{g}^{2})\|v\|_{2}^{2}.$
Taking the supremum over unit vectors proves the claim. ∎
Define the one-step contraction factor
$c_{x}(h):=\sqrt{1-2hm+h^{2}M_{g}^{2}}.$
For sufficiently small $h$, this factor is strictly less than one and governs depthwise propagation of perturbations. The next result translates this local bound into a global constraint on the forward analytic exponent.
Theorem 5.3 (Analytic stability of residual networks).
Consider the residual architecture (5.1) under assumption (5.2). Suppose the step size satisfies
$0<h<h_{\max},\qquad h_{\max}:=\frac{2m}{M_{g}^{2}}.$ (5.3)
Then $c_{x}(h)<1$ and the forward analytic exponent satisfies
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0.$
In particular, the forward component of the Learning Stability Profile is uniformly bounded in depth, and the hypotheses of Theorem 4.1 hold with a quadratic energy in $X_{k}$.
Sketch of proof.
By Lemma 5.2, each residual layer has spectral norm at most $c_{x}(h)$. The composition of $L$ layers therefore has forward Jacobian norm bounded by $c_{x}(h)^{L}$. If $h<h_{\max}$ then $1-2hm+h^{2}M_{g}^{2}<1$ and hence $c_{x}(h)<1$. The bound on $\alpha_{x}$ follows directly from Definition 3.1. Existence of a compatible learning energy then follows from Theorem 4.1. ∎
Condition (5.3) is a discrete stability restriction. Its structure is identical to a Courant–Friedrichs–Lewy type condition for explicit time stepping schemes applied to dissipative systems [Strikwerda2004]. The dimensionless quantity $hM_{g}$ measures the scale of each residual perturbation relative to intrinsic dissipation. If $h$ is too large, the contraction factor $c_{x}(h)$ exceeds one and the forward exponent becomes positive. The analytic stability framework therefore formalises the heuristic that residual blocks must be either sufficiently small in step size or strongly dissipative to ensure stable propagation [He2016ResNet, HaberRuthotto2017].
In many practical settings the parameters $\theta_{k}$ are updated by an inner gradient loop. Under the smoothness and convexity assumptions introduced in Section 6, such updates induce a parametric Lipschitz factor strictly smaller than one. Combining Theorem 5.3 with the stochastic gradient analysis yields negative forward and parametric exponents $(\alpha_{x},\alpha_{\theta})$, with $\alpha_{u}$ controlled by the noise level and learning rate. Residual networks satisfying the spectral and step-size conditions therefore lie in the analytically stable regime of Theorem 4.1.
Corollary 5.4 (CFL-type stability condition for residual networks).
Under the hypotheses of Theorem 5.3, the residual architecture is forward analytically stable whenever
$0<h<h_{\max},\qquad h_{\max}=\frac{2m}{M_{g}^{2}}.$
In this regime the forward analytic exponent satisfies
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0,$
so the representation dynamics are strictly contractive across depth.
6 Temporal stability of stochastic gradient methods
We consider stochastic gradient updates of the form
$\theta^{t+1}=\theta^{t}-\eta_{t}\,G(\theta^{t},u^{t}),\qquad t\geq 0,$ (6.1)
where $(\eta_{t})_{t\geq 0}$ is a step-size sequence and $(u^{t})_{t\geq 0}$ is an i.i.d. source of stochasticity.
Assumption 3 (Stochastic gradient regime).
The objective $\mathcal{L}$ is $\mu$–strongly convex and LL–smooth. The stochastic gradient oracle satisfies
$\mathbb{E}[G(\theta,u)\mid\theta]=\nabla\mathcal{L}(\theta),$
and there exist constants $\sigma_{0},\sigma_{1}\geq 0$ such that
$\mathbb{E}\!\left[\|G(\theta,u)-\nabla\mathcal{L}(\theta)\|^{2}\mid\theta\right]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|\theta-\theta^{\star}\|^{2}.$
Moreover, there exist constants $L_{G},L_{u}^{G}>0$ such that for all $\theta,\bar{\theta}\in\Theta$ and $u,\bar{u}\in\mathcal{U}$,
$\|G(\theta,u)-G(\bar{\theta},u)\|\leq L_{G}\|\theta-\bar{\theta}\|,\qquad\|G(\theta,u)-G(\theta,\bar{u})\|\leq L_{u}^{G}\|u-\bar{u}\|.$ (6.2)
6.1 Mean square Lyapunov recursion
We derive a mean square Lyapunov recursion for the parameter error. Define $e^{t}:=\theta^{t}-\theta^{\star}$ and decompose the stochastic gradient as
$G(\theta^{t},u^{t})=\nabla\mathcal{L}(\theta^{t})+\zeta^{t},\qquad\zeta^{t}:=G(\theta^{t},u^{t})-\nabla\mathcal{L}(\theta^{t}).$
By Assumption 3, $\mathbb{E}[\zeta^{t}\mid\theta^{t}]=0$ and
$\mathbb{E}\bigl[\|\zeta^{t}\|^{2}\mid\theta^{t}\bigr]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|e^{t}\|^{2}.$
The update (6.1) becomes
$e^{t+1}=e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})-\eta_{t}\zeta^{t}.$
Conditioning on $\theta^{t}$ and expanding yields
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]=\|e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})\|^{2}+\eta_{t}^{2}\mathbb{E}\bigl[\|\zeta^{t}\|^{2}\mid\theta^{t}\bigr].$
Strong convexity and smoothness of $\mathcal{L}$ imply
$\langle\nabla\mathcal{L}(\theta^{t}),e^{t}\rangle\geq\mu\|e^{t}\|^{2},\qquad\|\nabla\mathcal{L}(\theta^{t})\|^{2}\leq L^{2}\|e^{t}\|^{2}$
for all $\theta^{t}$ [Nesterov2004, Chapter 2]. Consequently,
$\|e^{t}-\eta_{t}\nabla\mathcal{L}(\theta^{t})\|^{2}\leq\bigl(1-2\eta_{t}\mu+\eta_{t}^{2}L^{2}\bigr)\|e^{t}\|^{2}.$
Combining the above estimates yields the Lyapunov recursion
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]\leq q_{t}\|e^{t}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2},$ (6.3)
where
$q_{t}:=1-2\eta_{t}\mu+\eta_{t}^{2}(L^{2}+\sigma_{1}^{2}).$
This inequality is the fundamental mean square energy estimate for the parameter energy
$\mathcal{E}(\theta):=\|\theta-\theta^{\star}\|^{2}.$
where the cross term vanishes since $\mathbb{E}[\zeta^{t}\mid\theta^{t}]=0$ by Assumption 3. The preceding computation can be summarized as the following lemma.
Lemma 6.1 (Mean square Lyapunov recursion for SGD).
Let Assumption 3 hold and let $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$. Then the SGD update (6.1) satisfies
$\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\mid\theta^{t}\bigr]\leq\mathcal{E}(\theta^{t})-\eta_{t}\bigl(2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\bigr)\mathcal{E}(\theta^{t})+\eta_{t}^{2}\sigma_{0}^{2}$
for all $t\geq 0$.
6.2 Temporal analytic stability for constant and decreasing step sizes
We now translate the Lyapunov recursion (6.3) into bounds on the parametric and temporal analytic exponents of the Learning Stability Profile.
Theorem 6.2 (Analytic temporal stability of SGD with constant step size).
Let Assumption 3 hold and suppose $\eta_{t}\equiv\eta$ is constant. Assume
$0<\eta<\eta_{\max},\qquad\eta_{\max}:=\frac{2\mu}{L^{2}+\sigma_{1}^{2}}.$ (6.4)
Then $q(\eta):=1-2\eta\mu+\eta^{2}(L^{2}+\sigma_{1}^{2})$ satisfies $0<q(\eta)<1$ and the following hold.
  1. 1.
    The quadratic energy $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$ obeys
    $\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\bigr]\leq q(\eta)\,\mathbb{E}\bigl[\mathcal{E}(\theta^{t})\bigr]+\eta^{2}\sigma_{0}^{2}$
    for all $t$. In particular, the mean square error converges to a stationary noise floor of order $\eta$.
  2. 2.
    The parametric analytic exponent satisfies
    $\alpha_{\theta}\leq\log q(\eta)<0.$
  3. 3.
    The temporal analytic exponent satisfies $\alpha_{u}\leq 0$. More precisely, there exists $C_{u}>0$, depending only on $(L_{G},L_{u}^{G},\mu,L)$ and on the chosen product norm on $\mathcal{M}$, such that
    $\|\mathsf{S}^{u}_{t}\|\leq C_{u}\,\eta\,L_{u}^{G}\qquad\text{for all }t.$
Sketch of proof.
Part (1) follows directly from (6.3) and the tower property of conditional expectation. Under (6.4), a standard linear recursion argument gives
$\mathbb{E}\bigl[\|e^{t}\|^{2}\bigr]\leq q(\eta)^{t}\|e^{0}\|^{2}+\frac{\eta^{2}\sigma_{0}^{2}}{1-q(\eta)}.$
For part (2), the Jacobian of the mean update map $\theta\mapsto\theta-\eta\nabla\mathcal{L}(\theta)$ has spectral radius at most $\sqrt{q(\eta)}$. This contraction rate governs the linearised dynamics, yielding $\alpha_{\theta}\leq\tfrac{1}{2}\log q(\eta)$ in the sense of Definition 3.1.
For part (3), differentiating the update map with respect to $u$ and using (6.2) gives an inhomogeneous linear system for perturbations driven by $\delta u^{t}$. Standard input–output bounds for linear time-varying systems [Khalil2002Nonlinear, Chapter 4] imply a uniform gain from $(\delta u^{0},\dots,\delta u^{t})$ to $\delta\theta^{t}$, yielding the stated bound and $\alpha_{u}\leq 0$. ∎
With constant step size, stability is not asymptotic in the deterministic sense, since the iterates fluctuate around $\theta^{\star}$ at a scale determined by $\eta$ and the noise variance. The parametric analytic exponent $\alpha_{\theta}$ nevertheless captures the exponential forgetting of initial conditions in mean square, while $\alpha_{u}$ quantifies the uniform sensitivity to perturbations in the stochastic update mechanism.
We now turn to the decreasing step size regime. In that case the iterates converge almost surely to $\theta^{\star}$ and both analytic exponents are nonpositive, in agreement with classical stochastic approximation theory [KushnerYin2003, Benaim1999].
Theorem 6.3 (Analytic temporal stability with decreasing step size).
Let Assumption 3 hold and suppose that the step sizes satisfy
$\eta_{t}>0,\qquad\sum_{t=0}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty.$
Then the following statements hold.
  1. 1.
    The sequence $(\theta^{t})$ converges almost surely to the unique minimiser $\theta^{\star}$. The energy $\mathcal{E}(\theta^{t})$ converges almost surely and is nonincreasing in expectation.
  2. 2.
    The parametric analytic exponent satisfies
    $\alpha_{\theta}\leq 0.$
  3. 3.
    The temporal analytic exponent satisfies
    $\alpha_{u}\leq 0.$
    More precisely, the temporal stability signatures $\sigma^{u}_{t}$ form a bounded sequence.
Sketch of proof.
The argument follows the classical ODE method for stochastic approximation [KushnerYin2003, Benaim1999]. Under Assumption 3, the mean dynamics associated with (6.1) track the gradient flow
$\dot{\theta}(s)=-\nabla\mathcal{L}(\theta(s)),$
which has a globally asymptotically stable equilibrium at $\theta^{\star}$ by strong convexity.
The step size conditions ensure that the stochastic perturbations are square summable while the mean drift is persistent. Standard almost supermartingale arguments (Robbins–Siegmund theorem; see [KushnerYin2003, Ch. 2]) imply that $\theta^{t}\to\theta^{\star}$ almost surely and that the Lyapunov energy $\mathcal{E}(\theta^{t})$ converges almost surely and decreases in expectation.
To obtain the analytic exponent bounds, note that convergence of $\theta^{t}$ together with the Jacobian bounds in (6.2) imply that the linearised parameter dynamics are asymptotically nonexpansive. Consequently, the parametric sensitivity operators remain uniformly bounded, yielding $\alpha_{\theta}\leq 0$.
Similarly, the Jacobian bound with respect to $u$ ensures that the linear response to perturbations of the update sequence remains uniformly bounded along the trajectory. Hence the temporal stability signatures $\sigma^{u}_{t}$ do not grow exponentially, and $\alpha_{u}\leq 0$. ∎
Theorems 6.2 and 6.3 place stochastic gradient methods squarely within the analytic stability framework developed in this paper. The quadratic energy $\mathcal{E}(\theta)$ acts as a Lyapunov function in expectation, while the learning rate conditions determine whether the parametric and temporal analytic exponents are strictly negative or merely nonpositive. In both regimes, the Learning Stability Profile provides a unified quantitative description of robustness to perturbations of initialisation and stochastic update mechanisms.
7 Variational Extension to Non-smooth Learning Systems
The analytic framework developed so far assumes that the layer maps $f_{k}$ and update maps $\mathcal{A}_{t}$ are continuously differentiable. Many architectures of practical interest violate this assumption. Common examples include ReLU and max pooling activations, hinge and $\ell_{1}$ losses, clipping and quantisation operators, and proximal or projected optimisation steps. This section extends the stability theory to such non-smooth learning systems.
The extension replaces classical Jacobians with Clarke generalized derivatives and smooth Lyapunov energies with variational energies that may be non differentiable. The resulting theory preserves the central equivalence established earlier: bounded stability signatures are equivalent to the existence of an energy that dissipates along learning trajectories. Thus the analytic exponents and stability laws remain valid in the non-smooth regime when derivatives are interpreted in the sense of Clarke [Clarke1990Optimization, Clarke1998Nonsmooth, RockafellarWets1998].
7.1 Locally Lipschitz learning flows and Clarke derivatives
We work in finite dimensional Euclidean spaces equipped with norms compatible with the product structure of the state space. The following assumption replaces Assumption 1.
Assumption 4 (Locally Lipschitz learning flow).
For each index $(k,t)$ the map
$\Phi_{k,t}\colon(x,\theta,u)\mapsto\bigl(f_{k}(x;\theta_{k}),\mathcal{A}_{t}(\theta,u)\bigr)$
is locally Lipschitz on $\mathcal{X}_{k}\times\Theta\times\mathcal{U}$. There exist constants
$L^{\mathrm{C}}_{\mathrm{fwd}},\qquad L^{\mathrm{C}}_{\mathrm{par}},\qquad L^{\mathrm{C}}_{\mathrm{upd}}$
such that for every $(x,\theta,u)$ and every Clarke generalized Jacobian
$V\in\partial_{C}\Phi_{k,t}(x,\theta,u),$
the blocks $V_{x}$, $V_{\theta}$, and $V_{u}$ satisfy
$\|V_{x}\|\leq L^{\mathrm{C}}_{\mathrm{fwd}},\qquad\|V_{\theta}\|\leq L^{\mathrm{C}}_{\mathrm{par}},\qquad\|V_{u}\|\leq L^{\mathrm{C}}_{\mathrm{upd}}.$
For a locally Lipschitz map $F\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$, the Clarke generalized Jacobian at $z$ is the compact convex set
$\partial_{C}F(z)=\mathrm{co}\Bigl\{\lim_{i\to\infty}JF(z_{i})\;\Big|\;z_{i}\to z,\,F\text{ differentiable at }z_{i}\Bigr\},$
where $JF(z_{i})$ denotes the classical Jacobian and $\mathrm{co}$ denotes convex hull [Clarke1990Optimization, Chapter 2]. Local Lipschitz continuity guarantees that $\partial_{C}F(z)$ is nonempty and upper semicontinuous in $z$ [Clarke1990Optimization, Clarke1998Nonsmooth].
Assumption 4 covers a wide class of non-smooth components used in learning systems. Piecewise affine activations such as ReLU and leaky ReLU are globally Lipschitz. Proximal operators of proper lower semicontinuous convex functionals are firmly nonexpansive and therefore Lipschitz with constant one [RockafellarWets1998]. Orthogonal and convex projections satisfy the same property [BauschkeCombettes2011]. Compositions of these maps with affine layers preserve local Lipschitz continuity and yield finite constants $L^{\mathrm{C}}_{\mathrm{fwd}}$, $L^{\mathrm{C}}_{\mathrm{par}}$, and $L^{\mathrm{C}}_{\mathrm{upd}}$.
Under Assumption 4, the learning dynamics are modelled as a discrete-time difference inclusion
$S^{t+1}_{k+1}\in\Phi_{k,t}(S^{t}_{k}),\qquad S^{t}_{k}=(X^{t}_{k},\theta^{t}),$
where the multivalued notation reflects possible non differentiability. Existence of trajectories follows from standard results for discrete-time systems with locally Lipschitz right-hand sides [Clarke1990Optimization, Chapter 2].
7.2 Generalized Learning Stability Profile and exponents
We now extend the Learning Stability Profile to the non-smooth setting. Sensitivities are measured using the largest block norms over all elements of the Clarke generalized Jacobian.
Definition 7.1 (Generalized Learning Stability Profile).
Under Assumption 4, the generalized stability signature at $(k,t)$ and state $S^{t}_{k}$ is the triple
$\mathrm{gLSP}(k,t;S^{t}_{k})=\bigl(\sigma^{x,C}_{k,t},\sigma^{\theta,C}_{k,t},\sigma^{u,C}_{k,t}\bigr),$
where
$\sigma^{x,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{x}\|,\quad\sigma^{\theta,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{\theta}\|,\quad\sigma^{u,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{u}\|.$
When $\Phi_{k,t}$ is differentiable at $S^{t}_{k}$, the Clarke generalized Jacobian reduces to the singleton containing the classical Jacobian, and the generalized signature coincides with the smooth Learning Stability Profile. In general, the generalized profile captures the worst-case local amplification of infinitesimal perturbations induced by any active linearisation of the non-smooth learning flow.
Definition 7.2 (Generalized analytic exponents).
Let $L$ and $T$ denote network depth and training horizon. The generalized analytic exponents are defined by
$\tilde{\alpha}_{x}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{x,C}_{k,t}\Bigr),$
$\tilde{\alpha}_{\theta}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{\theta,C}_{k,t}\Bigr),$
$\tilde{\alpha}_{u}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq t\leq T}\sigma^{u,C}_{k,t}\Bigr).$
Definition 7.3 (Generalized analytic stability).
The learning system is generalized analytically stable of order $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ if
$\tilde{\alpha}_{x}\leq\tilde{\beta}_{x},\qquad\tilde{\alpha}_{\theta}\leq\tilde{\beta}_{\theta},\qquad\tilde{\alpha}_{u}\leq\tilde{\beta}_{u},$
for some $\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u}\leq 0$.
These exponents extend the analytic stability exponents of Section 4 to locally Lipschitz learning flows. Negative values correspond to asymptotic contractivity of the generalized linearised dynamics, while nonpositive values correspond to uniform boundedness of generalized sensitivities.
7.3 Variational energies and dissipation
We now extend the energy side of the theory. Since non smooth energies may not admit classical gradients, we work with proper lower semicontinuous functionals and first order difference inequalities. This follows nonsmooth Lyapunov theory for differential inclusions, proximal algorithms, and stochastic approximation [Clarke1990Optimization, Benaim1999, BolteSabachTeboulle2014, DavisDrusvyatskiy2019].
Assumption 5 (Variational learning energy).
A variational learning energy is a proper lower semicontinuous functional
$\mathcal{E}\colon\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}\cup\{+\infty\}$
that is finite on all admissible learning states and satisfies the coercivity bounds
$c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}(X_{0},\dots,X_{L},\theta)\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr),$
for constants $0<c_{1}\leq c_{2}<\infty$.
Assumption 6 (Variational energy dissipation).
There exist a variational learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$ that is linear in $(\delta x,\delta\theta^{0},\delta u)$ at first order, such that for every training step $t$ and every admissible trajectory,
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u).$
Assumptions 5 and 6 are the non smooth analogues of Definition 3.3 and Assumption 2. Lower semicontinuity and coercivity ensure that $\mathcal{E}$ controls the norms of $X_{L}$ and $\theta$ even in the absence of differentiability, while the dissipation inequality encodes strict energy decay up to first order perturbations.
7.4 Fundamental Variational Stability Theorem
We now state the main result of the variational extension. It shows that the equivalence between bounded stability signatures and the existence of a dissipative energy persists in the non smooth regime.
Theorem 7.4 (Fundamental Variational Stability Theorem).
Let the flow maps $\Phi_{k,t}$ satisfy Assumption 4. Up to equivalence of norms on $\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$, the following statements are equivalent.
  1. 1.
    Generalized differential stability. There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all depths $L$, horizons $T$, and indices $0\leq k\leq L$, $0\leq t\leq T$,
    $\sigma^{x,C}_{k,t}\leq C_{x},\qquad\sigma^{\theta,C}_{k,t}\leq C_{\theta},\qquad\sigma^{u,C}_{k,t}\leq C_{u}.$
    Equivalently,
    $\tilde{\alpha}_{x}\leq 0,\qquad\tilde{\alpha}_{\theta}\leq 0,\qquad\tilde{\alpha}_{u}\leq 0.$
  2. 2.
    Variational energy dissipative stability. There exists a variational learning energy $\mathcal{E}$ satisfying Assumption 5 and a constant $\gamma>0$ such that the dissipation inequality in Assumption 6 holds for all learning trajectories.
Moreover, whenever these equivalent conditions hold, the generalized stability indices $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ in Definition 7.3 can be chosen explicitly as functions of $(L^{\mathrm{C}}_{\mathrm{fwd}},L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$ and $\gamma$.
Proof sketch.
The implication from (1) to (2) adapts the proof of Theorem 4.1. Uniform bounds on the generalized signatures imply incremental stability of the associated difference inclusion in a suitable product norm. Converse Lyapunov results for discrete time differential inclusions yield a regular Lyapunov function that decreases strictly along trajectories [Clarke1990Optimization, Clarke1998Nonsmooth]. By composing this function with the squared Euclidean norm of $(X_{L},\theta)$ one obtains a functional $\mathcal{E}$ satisfying Assumptions 5 and 6.
For the converse, coercivity of $\mathcal{E}$ controls $\|(X_{L},\theta)\|$ in terms of the energy. The dissipation inequality precludes unbounded growth of perturbations, since such growth would force the energy to increase along some subsequence. Selecting elements of $\partial_{C}\Phi_{k,t}(S^{t}_{k})$ and applying generalized chain rules [Clarke1990Optimization] yields uniform bounds on the linearised difference inclusion. These bounds imply uniform control of the generalized signatures and hence nonpositivity of the exponents. ∎
The theorem shows that analytic stability is preserved when passing from smooth to non smooth learning systems. The generalized exponents $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ retain the same interpretation as in the smooth case: negative values correspond to contractivity of the learning flow, while nonpositive values correspond to global boundedness of sensitivities with respect to inputs, parameters, and update perturbations.
7.5 Examples: piecewise linear networks and proximal flows
We briefly illustrate how the variational framework applies to two representative non-smooth systems. In each case we indicate (i) uniform bounds on the generalized stability signatures and (ii) existence of a variational energy satisfying a discrete dissipation inequality. Detailed derivations follow the same pattern as in Sections 5, 6, and 7 and are therefore omitted.
Piecewise linear networks.
Consider a feedforward architecture
$X_{k+1}=f_{k}(X_{k}):=\sigma(W_{k}X_{k}+b_{k}),\qquad k=0,\dots,L-1,$
where $\sigma$ is piecewise linear with slopes in $[0,1]$. Each $f_{k}$ is globally Lipschitz and differentiable almost everywhere. At differentiable points
$Jf_{k}(x)=D_{k}(x)W_{k},$
with $D_{k}(x)$ diagonal and entries in $[0,1]$. The Clarke generalized Jacobian satisfies
$\partial_{C}f_{k}(x)=\{DW_{k}\mid D\text{ diagonal},\ D_{ii}\in[0,1]\},$
so for any compatible operator norm
$\sup_{V\in\partial_{C}f_{k}(x)}\|V\|\leq\|W_{k}\|.$
If $\|W_{k}\|_{2}\leq\rho<1$ for all $k$, then Assumption 4 holds with $L^{\mathrm{C}}_{\mathrm{fwd}}\leq\rho$ and finite $(L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$. In particular,
$\sigma^{x,C}_{k,t}\leq\rho\quad\text{for all }(k,t),\qquad\tilde{\alpha}_{x}\leq\log\rho<0.$
For fixed parameters, the quadratic functional
$\mathcal{E}_{\mathrm{fwd}}(X_{0},\dots,X_{L},\theta):=\tfrac{1}{2}\|X_{L}\|^{2}$
is proper, lower semicontinuous, and coercive in $X_{L}$. Layerwise Lipschitz bounds yield a discrete Lyapunov inequality of the form Assumption 6 for a modified quadratic energy, with rate controlled by $1-\rho^{2}$. Hence the hypotheses of Theorem 7.4 are satisfied, and the generalized forward exponent can be read equivalently from Jacobian bounds or energy decay.
Proximal and projected flows.
Consider $\Phi(\theta)=\mathcal{L}(\theta)+R(\theta)$, where $\mathcal{L}$ has $L$-Lipschitz gradient and $R$ is proper, lower semicontinuous, and convex. The proximal gradient update
$\theta^{t+1}=\operatorname{prox}_{\eta R}\bigl(\theta^{t}-\eta\nabla\mathcal{L}(\theta^{t})\bigr)$
can be written as $T_{\eta}\circ S_{\eta}$, where $S_{\eta}(\theta)=\theta-\eta\nabla\mathcal{L}(\theta)$ and $T_{\eta}=\operatorname{prox}_{\eta R}$. For $0<\eta<2/L$, $S_{\eta}$ is nonexpansive and $T_{\eta}$ is firmly nonexpansive, hence $T_{\eta}\circ S_{\eta}$ is globally Lipschitz with constant at most one. Assumption 4 therefore holds with
$\sigma^{\theta,C}_{t}\leq 1\quad\text{for all }t,\qquad\tilde{\alpha}_{\theta}\leq 0.$
On the energy side, $\Phi$ itself is a variational learning energy. Under standard regularity conditions, the proximal gradient step satisfies
$\Phi(\theta^{t+1})-\Phi(\theta^{t})\leq-c_{\eta}\|\theta^{t+1}-\theta^{t}\|^{2},$
for some $c_{\eta}>0$, which is a special case of Assumption 6. The Fundamental Variational Stability Theorem 7.4 therefore applies to proximal and projected gradient flows, including the case where $R$ is the indicator of a closed convex set.
These examples show that architectural and optimisation-induced non-smoothness are handled within the same stability grammar by replacing classical Jacobians with Clarke derivatives and smooth energies with variational energies, while the generalized exponents and dissipation law retain the same form as in the smooth case.
8 Unified Interpretation
The results of this paper define an explicit stability grammar for learning systems. The (generalized) Learning Stability Profile characterises how perturbations in inputs, parameters, and update mechanisms propagate through the coupled representation–parameter flow. The Fundamental Analytic Stability Theorem and its variational counterpart show that uniform control of these sensitivities is equivalent to the existence of an energy functional that dissipates along learning trajectories, up to equivalence of norms on the joint state space.
8.1 Stability as a joint property of architecture and optimisation
In Sections 23, learning is modelled as a discrete-time dynamical system on a joint state space containing all layer representations and parameters. The Learning Stability Profile records operator norms of directional sensitivities derived from the Jacobian or generalized Jacobian of the coupled flow. The analytic and generalized exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ and $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ summarise the asymptotic growth or decay of these sensitivities along depth and training time.
The Fundamental Analytic Stability Theorem 4.1 and the Fundamental Variational Stability Theorem 7.4 show that, in both smooth and non-smooth regimes, nonpositive exponents are equivalent to the existence of an energy functional satisfying a discrete dissipation inequality. The choice between smooth and variational formulations is dictated by regularity: $C^{1}$ systems admit a Jacobian-based analysis, while locally Lipschitz systems require Clarke generalized derivatives and lower semicontinuous energies. On their overlap, the two formulations coincide, since the Clarke generalized Jacobian reduces to the classical Jacobian on a dense set and the associated exponents agree up to norm equivalence [Clarke1990Optimization, RockafellarWets1998].
This joint perspective links local analytic quantities, such as spectral norms, Lipschitz constants, and step sizes, to global dynamical behaviour through stability exponents and energies. Architecture and optimisation are treated symmetrically as components of a single coupled flow. The feedforward, residual, stochastic gradient, and variational stability laws derived in Sections 5, 6, and 7 are specialisations of this unified criterion.
8.2 Connection with classical stability notions
The analytic–variational framework unifies stability concepts across several communities. In numerical analysis and control, Lyapunov theory relates contractivity and negative exponents to strict Lyapunov functions [Khalil2002Nonlinear], while CFL-type conditions ensure stability of explicit schemes for dissipative systems [Strikwerda2004]. In optimisation, firm nonexpansiveness and averagedness yield contractive behaviour for proximal and splitting methods [BauschkeCombettes2011, BolteSabachTeboulle2014, DavisDrusvyatskiy2019]. In statistical learning, algorithmic stability bounds generalisation error via insensitivity to perturbations [Bousquet2002, Hardt2016Train, BottouCurtisNocedal2018].
Within the present framework, the Learning Stability Profile acts as a multi-directional Lyapunov exponent for the coupled learning flow. The analytic exponents reduce to spectral bounds for linear systems and to classical Lyapunov exponents in continuous-time limits, while the associated energy functionals generalise discrete Lyapunov functions to joint representation–parameter dynamics. Proximal and projected methods correspond to cases where the variational energy coincides with the objective, while algorithmic stability corresponds to nonpositive parametric and temporal exponents.
The feedforward and residual stability laws recover discrete counterparts of stability results for neural ordinary differential equations [HaberRuthotto2017, RuthottoHaber2020]. The CFL-type step size conditions correspond to admissible time steps for explicit discretisations of dissipative flows, while the stochastic gradient stability law recovers classical stochastic approximation conditions [KushnerYin2003, Benaim1999] in exponent form.
8.3 Design rules and trade offs
The explicit dependence of stability exponents on spectral norms, Lipschitz constants, dissipation rates, and learning rates yields concrete design rules. For feedforward networks, the forward exponent is controlled by layerwise spectral bounds, while for residual architectures it is governed by the product of the residual step size and the Jacobian bound of each block. The CFL-type constraint in Corollary 5.4 therefore acts as a design condition linking architectural scale and step size.
For stochastic gradient methods, the temporal stability law expresses the parametric exponent in terms of learning rate, curvature, and gradient noise. The effective contraction factor $q(\eta)$ in Theorem 6.2 refines deterministic stability criteria by incorporating stochastic effects, explaining the sensitivity of training to learning-rate schedules in high-variance regimes [BottouCurtisNocedal2018].
The framework also clarifies accuracy–stability trade offs. Impossibility results for certain expressive architectures [ColbrookAntilTownsend2021] correspond to regimes where the forward exponent must become positive as accuracy is pushed beyond a threshold. The analytic exponents thus provide a principled way to express stability constraints in terms of geometry and spectral scale rather than empirical performance alone.
8.4 Robustness, generalisation, and explanations
The analytic and generalized exponents connect stability to robustness and generalisation. Algorithmic stability theory links small generalisation gaps to insensitivity of learning procedures [Bousquet2002, Hardt2016Train]. In the present framework, bounded parametric and temporal exponents imply that perturbations in training data or stochasticity induce uniformly bounded changes in parameters and representations, as quantified by the Learning Stability Profile.
The same mechanism governs the stability of explanation methods. Explanation functionals based on gradients or saliency inherit their sensitivity from network Jacobians and from the stability of the training procedure [Fel2022]. When forward or parametric exponents are close to zero or positive, small perturbations may be amplified, leading to unstable explanations even when prediction error is low. Negative exponents instead imply quantitative robustness of both predictions and a broad class of gradient-based explanations. While interpretability also involves fidelity and domain alignment [Abusitta2024], the analytic–variational framework isolates stability as a structural property governed by architecture and optimisation.
8.5 Scope and limitations
The framework applies to finite-dimensional state spaces and discrete-time learning dynamics. The analytic results require global bounds on classical or generalized Jacobians, while the variational results require lower semicontinuous energies with quadratic coercivity along trajectories of interest. These assumptions cover many standard architectures and optimisers, including spectrally controlled feedforward networks, residual networks with bounded Jacobians, stochastic gradient methods under classical conditions, and proximal or projected flows.
They do not cover all regimes encountered in practice. Strongly adaptive optimisers, non-stationary data, heavy-tailed noise, or unbounded parameter growth may violate global bounds or admit only local or state-dependent energies [BottouCurtisNocedal2018]. The theory is therefore structural rather than fully quantitative: the derived inequalities relate exponents, regularity constants, and dissipation rates, but are not claimed to be sharp in high-dimensional settings.
The analytic and generalized exponents should thus be viewed as organising quantities capturing stability at the level of mechanisms and design rules, rather than as precise numerical predictors for specific large models. They unify Lyapunov stability, CFL-type conditions, proximal contraction, and algorithmic stability within a single analytic–variational framework.
9 Conclusion
This paper develops an analytic and variational framework for stability in deep learning systems. The central object is the Learning Stability Profile, which records the differential response of coupled representation, parameter, and update dynamics. The Fundamental Analytic Stability Theorem and its variational extension show that uniform bounds on these sensitivities are equivalent to the existence of Lyapunov-type energies dissipating along the joint learning flow, without requiring convexity, stationarity, or architectural homogeneity beyond stated regularity assumptions.
Within this framework, explicit stability laws are derived for standard model classes. Feedforward networks admit spectral conditions for forward contractivity, residual architectures satisfy discrete CFL-type constraints linking step size and Jacobian bounds, and stochastic gradient methods exhibit temporal stability governed by curvature, noise, and learning rate. In each case, stability is encoded by a small set of exponents and dissipation rates directly traceable to architectural and optimisation choices.
The framework extends naturally to non-smooth systems through Clarke generalized derivatives and variational energy methods. Piecewise linear networks, proximal and projected updates, and stochastic subgradient flows admit generalized stability signatures and Lyapunov energies satisfying the same dissipation law as in the smooth case. The resulting exponents capture contractivity even in the presence of kinks, projections, and non-smooth updates.
Several directions remain open. Continuous-time limits would connect the discrete exponents to Lyapunov exponents of neural ordinary differential equations [HaberRuthotto2017, RuthottoHaber2020]. Extensions to curved parameter or representation manifolds would introduce curvature-dependent stability indices and link the analysis to information geometry. Further work is needed to relate analytic and generalized exponents to empirical training behaviour in large-scale models and to develop practical diagnostics for estimating stability indices from data.
In conclusion, deep learning systems can be analysed as discrete dynamical systems governed by verifiable energy laws. The Learning Stability Profile provides a structural language that turns heuristic stabilisation rules into precise conditions on the joint learning flow, offering a foundation for principled architecture and optimisation design with explicit analytic stability guarantees.
Appendix A Detailed Proofs
This appendix presents proofs omitted from the main text.
A.1 Proofs for Section 3
Proof of Proposition 3.4.
Let $|\cdot|_{1}$ and $|\cdot|_{2}$ be two norms on the finite-dimensional space $\mathcal{M}$. By norm equivalence, there exist constants $m,M>0$ such that
$m|v|_{1}\leq|v|_{2}\leq M|v|_{1}\qquad\text{for all }v\in\mathcal{M}.$ (A.1)
For any linear operator $A:\mathcal{M}\to\mathcal{M}$, the induced operator norms satisfy the standard equivalence bounds
$|A|_{2,\mathrm{op}}\leq\tfrac{M}{m}|A|_{1,\mathrm{op}},\qquad|A|_{1,\mathrm{op}}\leq\tfrac{M}{m}|A|_{2,\mathrm{op}},$ (A.2)
see, e.g., [HornJohnson2012, Section 5.4].
Applying these inequalities to the directional sensitivity operators $\mathsf{S}^{x}_{k,t}$ yields constants $c=m/M$ and $C=M/m$ such that
$c\,\sigma^{x,1}_{k,t}\leq\sigma^{x,2}_{k,t}\leq C\,\sigma^{x,1}_{k,t}\quad\text{for all }k,t.$
Taking finite-horizon suprema and dividing by $L+T$ gives additive constants $\tfrac{1}{L+T}\log c$ and $\tfrac{1}{L+T}\log C$, which vanish as $L+T\to\infty$. Hence the limsup defining $\alpha_{x}$ is invariant under the choice of norm. The same argument applies verbatim to $\mathsf{S}^{\theta}_{k,t}$ and $\mathsf{S}^{u}_{k,t}$, proving the claim. ∎
Proof of Proposition 3.5.
We treat the forward component; the parameter and temporal components follow identically.
(i) Uniform boundedness implies $\alpha_{x}\leq 0$. If $\sigma^{x}_{k,t}\leq C_{x}$ for all $k,t$, then the finite-horizon supremum $M_{L,T}\leq C_{x}$, and hence
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log M_{L,T}\leq\lim_{L,T\to\infty}\frac{\log C_{x}}{L+T}=0.$
(ii) $\alpha_{x}<0$ implies exponential decay of finite-horizon suprema. If $\alpha_{x}<0$, then by definition of the limsup there exists $\varepsilon>0$ and $N\in\mathbb{N}$ such that
$\frac{1}{L+T}\log M_{L,T}\leq-\varepsilon\quad\text{whenever }L+T\geq N.$ (A.3)
Rearranging yields
$M_{L,T}\leq e^{-\varepsilon(L+T)}\quad\text{for all }L+T\geq N,$
which proves the claimed exponential decay. The same reasoning applies to $\alpha_{\theta}$ and $\alpha_{u}$. ∎
A.2 Proofs for section 4
Proof of Theorem 4.1.
We work on the finite-dimensional product space
$\mathcal{M}=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta$
equipped with an arbitrary norm $\|\cdot\|$ equivalent to a Euclidean norm. All constants below depend only on the bounds in Assumption 1 and on the chosen norm.
Denote by $S^{t}_{\cdot}\in\mathcal{M}$ the full learning state at training step $t$, and write the one-step learning map as
$S^{t+1}_{\cdot}=\Psi_{t}(S^{t}_{\cdot}),$
obtained by composing all layer maps and the parameter update at step $t$. Directional sensitivities with respect to $(\delta x,\delta\theta^{0},\delta u)$ are encoded by the Learning Stability Profile.
Step 1: (1) $\Rightarrow$ (2).
Assume item (1). Then there exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that
$\sigma^{x}_{k,t}\leq C_{x},\qquad\sigma^{\theta}_{k,t}\leq C_{\theta},\qquad\sigma^{u}_{k,t}\leq C_{u}\quad\text{for all }k,t.$
By definition of the directional sensitivities, this implies that the Fréchet derivative $D\Psi_{t}(S)$ is uniformly bounded in operator norm. Consequently, each $\Psi_{t}$ is globally Lipschitz on $\mathcal{M}$ with a constant $L_{\Psi}<\infty$ depending only on $(C_{x},C_{\theta},C_{u})$.
The assumption $\alpha_{x},\alpha_{\theta},\alpha_{u}\leq 0$ implies that, up to an equivalent choice of norm on $\mathcal{M}$, the difference dynamics
$\Delta S^{t+1}_{\cdot}=\Psi_{t}(S^{t}_{\cdot})-\Psi_{t}(\tilde{S}^{t}_{\cdot})$
are incrementally stable in the sense of discrete-time $\delta$–GAS [Angeli2002Incremental]. In particular, there exist constants $K\geq 1$ and $L_{\mathrm{inc}}\in(0,1]$ such that
$\|\Delta S^{t}_{\cdot}\|\leq KL_{\mathrm{inc}}^{t}\|\Delta S^{0}_{\cdot}\|\qquad\text{for all }t\geq 0.$ (A.4)
Strict negativity of the exponents yields $L_{\mathrm{inc}}<1$, while the marginal case corresponds to uniform boundedness.
By the discrete-time converse Lyapunov theorem for incrementally stable systems [Angeli2002Incremental, Theorem 3] (see also [Khalil2002Nonlinear, Chapter 13]), there exists a continuously differentiable incremental Lyapunov function
$V:\mathcal{M}\times\mathcal{M}\to\mathbb{R}_{+}$
satisfying quadratic bounds and a contraction inequality along the flow. Fixing a reference trajectory $\bar{S}^{t}_{\cdot}$ and defining
$\mathcal{E}(S):=V(S,\bar{S}^{t}_{\cdot}),$
yields an analytic learning energy that is coercive and satisfies
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2})$
for some $\gamma>0$ when $(\delta x,\delta\theta^{0},\delta u)=(0,0,0)$. When perturbations are present, standard smoothness arguments add a first-order remainder term linear in $(\delta x,\delta\theta^{0},\delta u)$, which is absorbed into $R$ in Assumption 2. This establishes item (2).
Step 2: (2) $\Rightarrow$ (1).
Assume item (2). Then there exist $\gamma>0$ and a $C^{1}$, coercive energy $\mathcal{E}$ satisfying
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2})+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$ (A.5)
with $R$ linear at first order.
For the unperturbed system, coercivity implies exponential decay of $\mathcal{E}(S^{t}_{\cdot})$ and hence global exponential stability of the learning dynamics in $(X_{L},\theta)$. Linearising the dynamics yields the variational system
$\delta S^{t+1}_{\cdot}=D\Psi_{t}(S^{t}_{\cdot})\,\delta S^{t}_{\cdot}.$
By discrete-time Lyapunov theory for linear time-varying systems [Khalil2002Nonlinear, Chapter 4], the inequality above implies uniform exponential bounds on the transition operators of the linearised system. Projecting onto the forward, parametric, and temporal directions yields uniform bounds on $\sigma^{x}_{k,t}$, $\sigma^{\theta}_{k,t}$, and $\sigma^{u}_{k,t}$.
By Definition 3.1, this implies $\alpha_{x},\alpha_{\theta},\alpha_{u}\leq 0$, establishing item (1). Moreover, the same Lyapunov argument yields explicit negative upper bounds on the stability indices $(\beta_{x},\beta_{\theta},\beta_{u})$ proportional to $\gamma$.
This completes the proof. ∎
Proof of Corollary 4.2.
Assume the hypotheses of Theorem 4.1.
(1) $\Rightarrow$ (2). If $\alpha_{x}\leq 0$, then by Definition 3.1 the forward signatures $(\sigma^{x}_{k,t})$ are uniformly bounded along the coupled trajectory. Hence there exists $C_{x}\geq 1$ such that
$\|\delta X^{t}_{k}\|\leq\sigma^{x}_{k,t}\|\delta x\|\leq C_{x}\|\delta x\|\qquad\text{for all }k,t.$
Applying Theorem 4.1 yields the existence of an analytic learning energy $\mathcal{E}$ satisfying the dissipation inequality in Assumption 2. Since the forward direction is treated symmetrically with the parametric and temporal directions, the associated remainder term is linear in $\delta x$ at first order. This establishes item (2).
(2) $\Rightarrow$ (1). Conversely, suppose that item (2) holds. Setting $(\delta\theta^{0},\delta u)=(0,0)$ and considering only input perturbations shows that the forward sensitivity operators are uniformly bounded. Proposition 3.5 then implies $\alpha_{x}\leq 0$, establishing item (1). ∎
Proof of Corollary 4.3.
Under the hypotheses of Theorem 4.1, the parametric and forward components of the Learning Stability Profile enter symmetrically.
(1) $\Rightarrow$ (2). If $\alpha_{\theta}\leq 0$, then the parametric signatures $\sigma^{\theta}_{k,t}$ are uniformly bounded. Hence there exists $C_{\theta}\geq 1$ such that for all initialization perturbations $\delta\theta^{0}$,
$\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{\theta}\|\delta\theta^{0}\|\qquad\text{for all }k,t.$
By Theorem 4.1, this uniform control implies the existence of an analytic learning energy $\mathcal{E}$ satisfying the dissipation inequality with a remainder term linear in $\delta\theta^{0}$, establishing item (2).
(2) $\Rightarrow$ (1). Conversely, if item (2) holds, the Lyapunov inequality applied to parametric perturbations yields uniform bounds on $\sigma^{\theta}_{k,t}$. Proposition 3.5 then implies $\alpha_{\theta}\leq 0$, proving item (1). ∎
Proof of Corollary 4.4.
The temporal component of the Learning Stability Profile is treated symmetrically with the forward and parametric components in Theorem 4.1.
(1) $\Rightarrow$ (2). If $\alpha_{u}\leq 0$, then the temporal signatures $\sigma^{u}_{k,t}$ are uniformly bounded. Hence there exists $C_{u}\geq 1$ such that for all admissible update perturbations $\delta u$,
$\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{u}\|\delta u\|\qquad\text{for all }k,t.$
Theorem 4.1 then guarantees the existence of an analytic learning energy $\mathcal{E}$ satisfying the dissipation inequality with a remainder linear in $\delta u$, proving item (2).
(2) $\Rightarrow$ (1). Conversely, if item (2) holds, the corresponding Lyapunov inequality implies uniform bounds on the temporal sensitivity operators $\mathsf{S}^{u}_{k,t}$. Proposition 3.5 then yields $\alpha_{u}\leq 0$, completing the proof. ∎
A.3 Proofs for section 5
Proof of Theorem 5.1.
Let the feedforward architecture satisfy the conditions in Section 5.1. Since $\sigma$ is 11–Lipschitz and differentiable almost everywhere, its Jacobian satisfies $\|D_{k}(x)\|_{2}\leq 1$ for all admissible $x$ (see, e.g., [RockafellarWets1998]).
Thus for each layer,
$Jf_{k}(x)=D_{k}(x)W_{k},\qquad\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}.$
By assumption $\|W_{k}\|_{2}\leq\rho<1$ uniformly in $k$. For the depth-$L$ network
$F_{\theta}=f_{L-1}\circ\cdots\circ f_{0},$
the chain rule yields
$JF_{\theta}(x)=Jf_{L-1}(X_{L-1})\cdots Jf_{0}(X_{0}),$
and submultiplicativity of the operator norm gives
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|Jf_{k}(X_{k})\|_{2}\leq\rho^{L}.$
Since in this architecture the forward sensitivity operator coincides with the Jacobian (Definition 3.1),
$\sigma^{x}_{L,t}=\|\mathsf{S}^{x}_{L,t}\|\leq\rho^{L}.$
Hence
$\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\leq\rho^{L}.$
By Definition 3.1,
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\!\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)=\log\rho<0.$
Negativity of the forward exponent implies analytic forward stability. By Theorem 4.1, the uniform spectral contraction ensures the existence of an analytic learning energy $\mathcal{E}$ satisfying the dissipation law in Assumption 2. Since the architecture is purely feedforward, $\mathcal{E}$ may be chosen to depend only on $(X_{L},\theta)$. ∎
Proof of Lemma 5.2.
Consider a residual layer
$X_{k+1}=X_{k}+h\,g_{k}(X_{k};\theta_{k}),$
with Jacobian $G_{k}(x)=\partial_{x}g_{k}(x;\theta_{k})$. The Jacobian of the residual map is
$J_{k}=I+hG_{k}.$
For any unit vector $v$,
$\|J_{k}v\|_{2}^{2}=\|v\|_{2}^{2}+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
By the dissipativity and boundedness assumptions (5.2),
$v^{\top}G_{k}v\leq-m\|v\|_{2}^{2},\qquad\|G_{k}v\|_{2}\leq M_{g}\|v\|_{2},$
whence
$\|J_{k}v\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2}.$
Taking the supremum over all unit vectors yields
$\|J_{k}\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2},$
as claimed. ∎
Proof of Theorem 5.3.
Lemma 5.2 yields the uniform bound
$\biggl\|\frac{\partial X_{k+1}}{\partial X_{k}}\biggr\|_{2}\leq c_{x}(h):=\sqrt{1-2hm+h^{2}M_{g}^{2}}.$
For the depth-$L$ residual composition
$F_{\theta}=f_{L-1}\circ\cdots\circ f_{0},\quad f_{k}(x)=x+hg_{k}(x;\theta_{k}),$
the chain rule gives
$\|JF_{\theta}(x)\|_{2}\leq c_{x}(h)^{L}.$
If $0<h<h_{\max}=2m/M_{g}^{2}$, then $c_{x}(h)<1$. Hence the forward signatures satisfy
$\sigma^{x}_{k,t}\leq c_{x}(h)^{L},$
uniformly in $(k,t)$. By Definition 3.1,
$\alpha_{x}=\log c_{x}(h)=\frac{1}{2}\log(1-2hm+h^{2}M_{g}^{2})<0.$
Thus the representation dynamics are exponentially contractive across depth. By Theorem 4.1, this implies the existence of a quadratic analytic learning energy satisfying the energy–dissipation inequality in Assumption 2. ∎
Proof of Corollary 5.4.
Under $0<h<h_{\max}$ we have $c_{x}(h)<1$ and hence
$\alpha_{x}=\frac{1}{2}\log(1-2hm+h^{2}M_{g}^{2})<0$
by Theorem 5.3. The forward sensitivity signatures are therefore uniformly contractive across depth, and Theorem 4.1 yields analytic forward stability of the residual architecture in the sense of Definition 3.2. ∎
A.4 Proofs for section 6
We work throughout under Assumption 3 and with the stochastic gradient recursion
$\theta^{t+1}=\theta^{t}-\eta_{t}G(\theta^{t},u^{t}),\qquad t\geq 0,$ (A.6)
where $(u^{t})_{t\geq 0}$ is an i.i.d. sequence on $\mathcal{U}$, independent of $\theta^{0}$.
Temporal stability: constant step size
We now give a complete proof of Theorem 6.2. Throughout this subsection we assume $\eta_{t}\equiv\eta$ and use the shorthand
$q(\eta):=1-2\eta\mu+\eta^{2}(L^{2}+\sigma_{1}^{2}).$
Theorem A.1 (Analytic temporal stability of SGD with constant step size).
Let Assumption 3 hold and let $\eta_{t}\equiv\eta$ be constant. Suppose that
$0<\eta<\eta_{\max},\qquad\eta_{\max}:=\frac{2\mu}{L^{2}+\sigma_{1}^{2}}.$ (A.7)
Then $q(\eta)\in(0,1)$ and the following statements hold.
  1. 1.
    The quadratic energy $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$ satisfies
    $\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\bigr]\;\leq\;q(\eta)\,\mathbb{E}\bigl[\mathcal{E}(\theta^{t})\bigr]+\eta^{2}\sigma_{0}^{2},\qquad t\geq 0.$ (A.8)
    In particular,
    $\mathbb{E}\bigl[\mathcal{E}(\theta^{t})\bigr]\;\leq\;q(\eta)^{t}\,\mathcal{E}(\theta^{0})+\frac{\eta^{2}\sigma_{0}^{2}}{1-q(\eta)},\qquad t\geq 0.$ (A.9)
  2. 2.
    The parametric analytic exponent associated with the Lyapunov energy $\mathcal{E}$ satisfies
    $\alpha_{\theta}\;\leq\;\log q(\eta)\;<\;0.$ (A.10)
  3. 3.
    The temporal analytic exponent satisfies $\alpha_{u}\leq 0$. More precisely, there exists a constant $C_{u}>0$, depending only on $(L_{G},L_{u}^{G},\mu,L)$ and the choice of equivalent product norm on $\mathcal{M}$, such that the temporal sensitivity operators obey
    $\|\mathsf{S}^{u}_{t}\|\;\leq\;C_{u}\eta L_{u}^{G}\qquad\text{for all }t\geq 0.$ (A.11)
Proof.
Step 1: Energy recursion and stationary bound. Taking expectations in Lemma 6.1 (with $\eta_{t}\equiv\eta$) yields (A.8). Unrolling the scalar inequality $a_{t+1}\leq q(\eta)a_{t}+b$ with $a_{t}=\mathbb{E}[\mathcal{E}(\theta^{t})]$ and $b=\eta^{2}\sigma_{0}^{2}$ gives (A.9) whenever $q(\eta)\in(0,1)$.
It remains to verify $q(\eta)\in(0,1)$ under (A.7). The bound $q(\eta)<1$ is immediate from $q(\eta)=1-\eta(2\mu-\eta(L^{2}+\sigma_{1}^{2}))$ and $\eta<2\mu/(L^{2}+\sigma_{1}^{2})$. For positivity, note that $L\geq\mu$ under Assumption 3 (since $\mu I\preceq\nabla^{2}\mathcal{L}\preceq LI$), hence $L^{2}+\sigma_{1}^{2}\geq\mu^{2}$ and
$q(\eta)=(L^{2}+\sigma_{1}^{2})\Bigl(\eta-\frac{\mu}{L^{2}+\sigma_{1}^{2}}\Bigr)^{2}+\Bigl(1-\frac{\mu^{2}}{L^{2}+\sigma_{1}^{2}}\Bigr)\geq 0,$
with strict positivity for $\eta\in(0,\eta_{\max})$ unless $\mu=L$ and $\sigma_{1}=0$ and $\eta=\mu/(L^{2}+\sigma_{1}^{2})$, a degenerate boundary case excluded by strict inequality in (A.7). Thus $q(\eta)\in(0,1)$.
Step 2: Parametric analytic exponent. Consider two trajectories $\theta^{t},\bar{\theta}^{t}$ driven by the same noise $(u^{t})$ with different initializations. Writing $d^{t}:=\theta^{t}-\bar{\theta}^{t}$ and applying the same mean–square Lyapunov argument as Lemma 6.1 to the coupled recursion (standard for contractive stochastic approximation; see [KushnerYin2003, Ch. 2, Ch. 4]), one obtains
$\mathbb{E}\|d^{t+1}\|^{2}\leq q(\eta)\,\mathbb{E}\|d^{t}\|^{2},\qquad t\geq 0,$
and hence $\mathbb{E}\|d^{t}\|^{2}\leq q(\eta)^{t}\|d^{0}\|^{2}$. Interpreting $d^{t}$ as the image of the Fréchet derivative of the flow map $\theta^{0}\mapsto\theta^{t}$ along direction $d^{0}$ yields the induced sensitivity decay $\|\mathsf{S}^{\theta}_{k,t}\|\lesssim q(\eta)^{t/2}$ (up to equivalence of norms on $\mathcal{M}$). Therefore
$\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{\theta}_{k,t}\leq C\,q(\eta)^{T/2},$
and Definition 3.1 gives $\alpha_{\theta}\leq\tfrac{1}{2}\log q(\eta)\leq\log q(\eta)<0$, proving (A.10).
Step 3: Temporal analytic exponent and (A.11). Let $u$ and $\bar{u}$ differ only at a single index $s$. With the same initialization, define $\delta\theta^{t}:=\theta^{t}-\bar{\theta}^{t}$. Then $\delta\theta^{s}=0$ and
$\delta\theta^{s+1}=-\eta\bigl(G(\theta^{s},u^{s})-G(\theta^{s},\bar{u}^{s})\bigr),\qquad\|\delta\theta^{s+1}\|\leq\eta L_{u}^{G}\|u^{s}-\bar{u}^{s}\|.$
For $t\geq s+1$,
$\delta\theta^{t+1}=\delta\theta^{t}-\eta\bigl(G(\theta^{t},u^{t})-G(\bar{\theta}^{t},u^{t})\bigr),$
and the Lipschitz bound $\|G(\theta,u)-G(\bar{\theta},u)\|\leq L_{G}\|\theta-\bar{\theta}\|$ yields $\|\delta\theta^{t+1}\|\leq(1+\eta L_{G})\|\delta\theta^{t}\|$. Combining these one-step bounds and summing the effects over $s$ (with respect to an equivalent product norm on $\mathcal{U}^{\mathbb{N}}$ adapted to the stability profile) gives the uniform operator bound (A.11) (this is the standard discrete-time input-to-state stability estimate; see, e.g., [Khalil2002Nonlinear, Ch. 4]). In particular $\sigma^{u}_{t}=\|\mathsf{S}^{u}_{t}\|$ is bounded in $t$, hence $\alpha_{u}\leq 0$ by Definition 3.1. ∎
Temporal stability: decreasing step size
We now prove Theorem 6.3 in a form adapted to the analytic stability framework.
Theorem A.2 (Analytic temporal stability with decreasing step size).
Let Assumption 3 hold and suppose that the stepsizes satisfy
$\eta_{t}>0,\qquad\sum_{t=0}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty.$ (A.12)
Then:
  1. 1.
    The sequence $(\theta^{t})_{t\geq 0}$ converges almost surely to the unique minimiser $\theta^{\star}$ of $\mathcal{L}$. The Lyapunov energy $\mathcal{E}(\theta^{t})=\|\theta^{t}-\theta^{\star}\|^{2}$ converges almost surely and is nonincreasing in expectation in the sense that
    $\sum_{t=0}^{\infty}\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})-\mathcal{E}(\theta^{t})\bigr]>-\infty.$ (A.13)
  2. 2.
    The parametric analytic exponent satisfies $\alpha_{\theta}\leq 0$.
  3. 3.
    The temporal analytic exponent satisfies $\alpha_{u}\leq 0$. More precisely, the temporal signatures $\sigma^{u}_{t}=\|\mathsf{S}^{u}_{t}\|$ form a bounded sequence in $t$.
Proof.
Step 1: Almost sure convergence and Lyapunov property. Lemma 6.1 gives
$\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\mid\theta^{t}\bigr]\leq\mathcal{E}(\theta^{t})-\eta_{t}\Bigl(2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\Bigr)\mathcal{E}(\theta^{t})+\eta_{t}^{2}\sigma_{0}^{2}.$
Since $\eta_{t}\to 0$ and $\sum_{t}\eta_{t}^{2}<\infty$, for all large $t$ we have $2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\geq\mu$, hence
$\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\mid\theta^{t}\bigr]\leq\mathcal{E}(\theta^{t})-\mu\eta_{t}\mathcal{E}(\theta^{t})+\eta_{t}^{2}\sigma_{0}^{2}.$
Taking expectations and summing over $t$ yields a supermartingale-type estimate implying $\sum_{t}\eta_{t}\,\mathbb{E}[\mathcal{E}(\theta^{t})]<\infty$ and (A.13). By the Robbins–Siegmund theorem ([Benaim1999, Theorem 1]) the process $\mathcal{E}(\theta^{t})$ converges almost surely and $\sum_{t}\eta_{t}\mathcal{E}(\theta^{t})<\infty$ a.s. Since $\sum_{t}\eta_{t}=\infty$, this forces $\mathcal{E}(\theta^{t})\to 0$ a.s., and strong convexity then implies $\theta^{t}\to\theta^{\star}$ a.s. (standard stochastic approximation; see [KushnerYin2003, Theorem 2.1, Ch. 2]).
Step 2: Parametric analytic exponent. Let $\theta^{t},\bar{\theta}^{t}$ be driven by the same noise with different initializations and set $d^{t}:=\theta^{t}-\bar{\theta}^{t}$. Applying the same mean–square Lyapunov calculation to the coupled difference recursion (as in Step 2 of Theorem A.1, but with $\eta_{t}$ varying) yields an inequality of the form
$\mathbb{E}\|d^{t+1}\|^{2}\leq\bigl(1-c\eta_{t}+C\eta_{t}^{2}\bigr)\,\mathbb{E}\|d^{t}\|^{2}\qquad(t\ \text{large}),$
for constants $c>0$, $C\geq 0$ depending only on Assumption 3 (see [KushnerYin2003, Ch. 4] for this standard stability step). Since $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$, the product of factors $\prod_{s\leq t}(1-c\eta_{s}+C\eta_{s}^{2})$ is finite and in fact tends to 0. In particular $\sup_{t}\mathbb{E}\|d^{t}\|^{2}\leq C_{\theta}\|d^{0}\|^{2}$, so the induced operator norms of the parametric sensitivity operators are uniformly bounded (up to equivalence of norms on $\mathcal{M}$). Hence
$\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{\theta}_{k,t}\leq C_{\theta}^{\prime},$
with $C_{\theta}^{\prime}$ independent of $(L,T)$, and Definition 3.1 gives $\alpha_{\theta}\leq 0$.
Step 3: Temporal analytic exponent. Let $u,\bar{u}$ differ at finitely many indices and define $\delta\theta^{t}:=\theta^{t}-\bar{\theta}^{t}$ for trajectories with the same initialization. The one-step decomposition gives
$\delta\theta^{t+1}=\delta\theta^{t}-\eta_{t}\bigl(G(\theta^{t},u^{t})-G(\bar{\theta}^{t},\bar{u}^{t})\bigr)=\delta\theta^{t}-\eta_{t}\bigl(G(\theta^{t},u^{t})-G(\bar{\theta}^{t},u^{t})\bigr)-\eta_{t}\bigl(G(\bar{\theta}^{t},u^{t})-G(\bar{\theta}^{t},\bar{u}^{t})\bigr).$
Using $\|G(\theta,u)-G(\bar{\theta},u)\|\leq L_{G}\|\theta-\bar{\theta}\|$ and $\|G(\theta,u)-G(\theta,\bar{u})\|\leq L_{u}^{G}\|u-\bar{u}\|$ gives the ISS-type bound
$\|\delta\theta^{t+1}\|\leq(1+\eta_{t}L_{G})\|\delta\theta^{t}\|+\eta_{t}L_{u}^{G}\|u^{t}-\bar{u}^{t}\|.$
Iterating and using $\eta_{t}\to 0$ with $\sum_{t}\eta_{t}^{2}<\infty$ (together with an equivalent product norm on $\mathcal{U}^{\mathbb{N}}$ compatible with the analytic stability profile) yields a uniform operator bound
$\|\mathsf{S}^{u}_{t}\|\leq C_{u}\eta_{\max}L_{u}^{G}\qquad(t\geq 0),$
for some finite $C_{u}>0$ depending only on the regime constants and norm choice (standard discrete-time ISS reasoning; see [Khalil2002Nonlinear, Ch. 4]). Thus $\sigma^{u}_{t}$ is bounded in $t$ and Definition 3.1 gives $\alpha_{u}\leq 0$. ∎
A.5 Proofs for section 7
Lemma A.3 (Smooth case as a special case of the generalized LSP).
Assume that, for a fixed pair $(k,t)$ and state $S^{t}_{k}$, the map $\Phi_{k,t}$ is differentiable at $S^{t}_{k}$ in the classical sense. Let $J\Phi_{k,t}(S^{t}_{k})$ denote the Jacobian and let $\sigma^{x}_{k,t},\sigma^{\theta}_{k,t},\sigma^{u}_{k,t}$ be the (smooth) stability signatures from Definition 3.1. Let $\sigma^{x,C}_{k,t},\sigma^{\theta,C}_{k,t},\sigma^{u,C}_{k,t}$ denote the generalized signatures from Definition 7.1. Then
$\partial_{C}\Phi_{k,t}(S^{t}_{k})=\{J\Phi_{k,t}(S^{t}_{k})\}\quad\text{and}\quad\sigma^{x,C}_{k,t}=\sigma^{x}_{k,t},\quad\sigma^{\theta,C}_{k,t}=\sigma^{\theta}_{k,t},\quad\sigma^{u,C}_{k,t}=\sigma^{u}_{k,t}.$
In particular the generalized exponents $\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u}$ coincide with the analytic exponents $\alpha_{x},\alpha_{\theta},\alpha_{u}$ whenever all $\Phi_{k,t}$ are smooth.
Proof.
By the standard reduction of the Clarke generalized Jacobian to the classical Jacobian at points of differentiability [Clarke1990Optimization, Prop. 2.6.2], $\partial_{C}\Phi_{k,t}(S^{t}_{k})=\{J\Phi_{k,t}(S^{t}_{k})\}$. Since the block operator norms in Definitions 3.1 and 7.1 agree when the generalized Jacobian is a singleton, the generalized and smooth signatures coincide. The equality of exponents follows immediately from Definitions 3.1 and 7.2. ∎
Remark A.4 (Two variational proof frameworks).
The Fundamental Variational Stability Theorem admits two complete proof frameworks.
  1. 1.
    Option A: Clarke incremental stability route. Use the Clarke generalised Jacobian, uniform bounds on the generalized signatures, and incremental stability theory for discrete-time differential inclusions to construct a (possibly nonsmooth) Lyapunov function decreasing along trajectories; see [Clarke1990Optimization, Clarke1998Nonsmooth, RockafellarWets1998]. This yields a variational energy $\mathcal{E}$ and the dissipation inequality of Assumption 6.
  2. 2.
    Option B: variational descent route. Work directly with a proper l.s.c. energy and apply nonsmooth chain rules / descent inequalities for difference inclusions, as in modern nonsmooth descent theory [BolteSabachTeboulle2014, DavisDrusvyatskiy2019]. Under bounded generalized signatures the learning flow yields strict energy decrease up to first-order perturbation terms, again producing Assumption 6.
Both routes apply under Assumption 4 and yield the same conclusions for generalized exponents and dissipative variational energies.
Theorem A.5 (Fundamental Variational Stability Theorem).
Let the flow maps $\Phi_{k,t}$ satisfy Assumption 4. Up to equivalence of norms on $\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$ the following statements are equivalent.
  1. 1.
    Generalized differential stability holds. There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all depths $L$, horizons $T$, and indices $0\leq k\leq L$, $0\leq t\leq T$,
    $\sigma^{x,C}_{k,t}\leq C_{x},\qquad\sigma^{\theta,C}_{k,t}\leq C_{\theta},\qquad\sigma^{u,C}_{k,t}\leq C_{u}.$
    Equivalently,
    $\tilde{\alpha}_{x}\leq 0,\qquad\tilde{\alpha}_{\theta}\leq 0,\qquad\tilde{\alpha}_{u}\leq 0.$
  2. 2.
    Variational energy dissipative stability holds. There exists a variational learning energy $\mathcal{E}$ satisfying Assumption 5 and a constant $\gamma>0$ such that the dissipation inequality in Assumption 6 holds for all learning trajectories.
Moreover, whenever these equivalent conditions hold, one can choose generalized stability indices $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ in Definition 7.3 as explicit functions of $(L^{\mathrm{C}}_{\mathrm{fwd}},L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$ and $\gamma$.
(Option A: Clarke incremental stability).
Work on the finite-dimensional product space
$\mathcal{M}:=\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$
with any Euclidean norm $\|\cdot\|$ (all such norms are equivalent).
Step 1: (1) $\Rightarrow$ incremental exponential stability. By Assumption 4, each $\Phi_{k,t}$ is locally Lipschitz, hence Clarke regular with nonempty compact upper semicontinuous generalized Jacobian $\partial_{C}\Phi_{k,t}$ [Clarke1990Optimization, Thm. 2.6.6]. The block bounds in Assumption 4 imply that for every $V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})$,
$\|V\|\leq L_{\mathrm{tot}}:=L^{\mathrm{C}}_{\mathrm{fwd}}+L^{\mathrm{C}}_{\mathrm{par}}+L^{\mathrm{C}}_{\mathrm{upd}},$
after passing to an equivalent product norm if needed.
For any two solutions $\{S^{t}_{k}\}$ and $\{\bar{S}^{t}_{k}\}$ driven by the same update indices, the Clarke mean value inequality [Clarke1990Optimization, Thm. 2.6.5] gives the one-step estimate
$\|S^{t+1}_{k+1}-\bar{S}^{t+1}_{k+1}\|\leq L_{\mathrm{tot}}\|S^{t}_{k}-\bar{S}^{t}_{k}\|.$
Iterating along the $(k,t)$-grid yields finite-product bounds in terms of products of generalized Jacobians. Under item (1), the generalized signatures are uniformly bounded; equivalently, the set of admissible generalized Jacobians has bounded joint spectral radius. Therefore one may select an equivalent norm $\|\cdot\|_{\ast}$ for which the induced one-step Lipschitz constant is $\lambda\in(0,1)$ (standard joint-spectral-radius norm construction; see [RockafellarWets1998, Thm. 5.7.11]). Hence
$\|S^{t+1}_{k+1}-\bar{S}^{t+1}_{k+1}\|_{\ast}\leq\lambda\|S^{t}_{k}-\bar{S}^{t}_{k}\|_{\ast}\quad\text{for all }k,t,$ (A.14)
i.e. the inclusion is incrementally exponentially stable (cf. [Angeli2002Incremental]).
Step 2: Incremental stability $\Rightarrow$ Lyapunov function for the inclusion. By the converse Lyapunov theorem for incrementally exponentially stable difference inclusions, there exists a proper l.s.c. locally Lipschitz Lyapunov function $V:\mathcal{M}\to\mathbb{R}_{+}$ and constants $a_{1},a_{2},a_{3}>0$ such that [Clarke1998Nonsmooth, Thm. 3.2, Thm. 5.1]
$a_{1}\|z\|_{\ast}^{2}\leq V(z)\leq a_{2}\|z\|_{\ast}^{2},\qquad V(S^{t+1}_{k+1})-V(S^{t}_{k})\leq-a_{3}\|S^{t}_{k}\|_{\ast}^{2},$
along every solution.
Step 3: Construct $\mathcal{E}$ and obtain Assumption 6. Define the learning energy
$\mathcal{E}(X_{0},\dots,X_{L},\theta):=V(X_{0},\dots,X_{L},\theta).$
The quadratic comparison bounds and norm equivalence give the coercivity requirement in Assumption 5 (possibly after adjusting constants and restricting to the $(X_{L},\theta)$-coordinates up to equivalence). Lower semicontinuity and local Lipschitzness follow from the construction [Clarke1990Optimization, Sec. 2.5]. The decay bound implies
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\|S^{t}_{\cdot}\|_{\ast}^{2}$
for some $\gamma>0$. Expressing $\|S^{t}_{\cdot}\|_{\ast}^{2}$ in terms of $\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}$ via norm equivalence yields the dissipation inequality in Assumption 6 for unperturbed dynamics. For primitive perturbations $(\delta x,\delta\theta^{0},\delta u)$, apply the same inequality to perturbed vs. unperturbed trajectories; local Lipschitz regularity yields a first-order remainder term $R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u)$ bounded linearly in the perturbations, giving Assumption 6.
Step 4: (2) $\Rightarrow$ (1). Assume item (2) holds with $\mathcal{E}$ and $\gamma>0$. Coercivity gives $c_{1}\|S\|^{2}\leq\mathcal{E}(S)\leq c_{2}\|S\|^{2}$ and norm equivalence $C_{\mathrm{eq}}^{-1}\|S\|\leq\|S\|_{\ast}\leq C_{\mathrm{eq}}\|S\|$ for some equivalent $\|\cdot\|_{\ast}$. If, say, $\sigma^{x,C}_{k,t}$ were unbounded, one could choose $(k_{n},t_{n})$, $S^{t_{n}}_{k_{n}}$ and $V_{n}\in\partial_{C}\Phi_{k_{n},t_{n}}(S^{t_{n}}_{k_{n}})$ with $\|V_{n,x}\|\to\infty$, and directions $h_{n}$ with $\|h_{n}\|_{\ast}=1$ and $\|V_{n}h_{n}\|_{\ast}\geq\tfrac{1}{2}\|V_{n,x}\|$. Perturbing the state by $\varepsilon h_{n}$ and using the Clarke chain rule / directional derivative representation [Clarke1990Optimization, Sec. 2.2] forces an arbitrarily large first-order variation in $\mathcal{E}$, contradicting the uniform dissipation inequality in Assumption 6 (since $\partial_{C}\mathcal{E}$ is bounded on bounded sets for locally Lipschitz $\mathcal{E}$; see [Clarke1990Optimization, Prop. 2.1.2]). The same contradiction argument applies to the parameter and update blocks. Hence there exist $C_{x},C_{\theta},C_{u}$ such that for all $k,t$,
$\sigma^{x,C}_{k,t}\leq C_{x},\quad\sigma^{\theta,C}_{k,t}\leq C_{\theta},\quad\sigma^{u,C}_{k,t}\leq C_{u},$
and therefore $\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u}\leq 0$.
Step 5: Stability indices. Given $(C_{x},C_{\theta},C_{u})$ and $\gamma$, standard joint-spectral-radius and norm-equivalence bounds for products of admissible generalized Jacobians yield explicit nonpositive (and in the strictly dissipative case, negative) upper bounds $\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u}$ depending only on $(L^{\mathrm{C}}_{\mathrm{fwd}},L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$ and $\gamma$; see, e.g., [Clarke1998Nonsmooth, Sec. 3.3] and [Khalil2002Nonlinear, Ch. 2]. This completes the Option A proof. ∎
(Option B: variational descent route).
We sketch a proof that works directly with descent inequalities for a variational energy, without explicitly invoking incremental stability. Work again on $\mathcal{M}$ with any Euclidean norm.
Step 1: (1) $\Rightarrow$ (2) via converse Lyapunov / descent construction. Assume item (1). For fixed $L$, write the time-$t$ one-step map on $\mathcal{M}$ as $T_{t}(S):=\Phi_{\cdot,t}(S)$. Bounded generalized signatures imply a uniform Clarke-Lipschitz bound $\|T_{t}(S)-T_{t}(\bar{S})\|\leq L_{\max}\|S-\bar{S}\|$ for some finite $L_{\max}\geq 1$ by the Clarke mean value theorem [Clarke1990Optimization, Thm. 2.6.5]. Choose $\lambda>0$ so that $\lambda L_{\max}^{2}<1$ and define the standard converse-Lyapunov energy
$\mathcal{E}(S):=\sup_{n\geq 0}\lambda^{-n}\|T_{n-1}\circ\cdots\circ T_{0}(S)\|^{2}.$ (A.15)
By the classical discrete-time converse Lyapunov construction (adapted to time-varying Lipschitz maps) the supremum is finite, and $\mathcal{E}$ is proper, l.s.c., and coercive with quadratic comparison bounds [Khalil2002Nonlinear, Thm. 4.1]; the l.s.c. property follows from stability of suprema of continuous functions and finiteness on trajectories.
Moreover, by construction one has a one-step decay $\mathcal{E}(S^{t+1}_{\cdot})\leq\lambda\,\mathcal{E}(S^{t}_{\cdot})$ for unperturbed dynamics, and for perturbed dynamics the nonsmooth chain rule plus the standard descent inequality for l.s.c. energies yields
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-(1-\lambda)\mathcal{E}(S^{t}_{\cdot})+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$
with $R$ linear in the primitive perturbations at first order (see, e.g., [Clarke1990Optimization, Prop. 2.1.2] and [BolteSabachTeboulle2014, Prop. 3.1]). Combining with coercivity and norm equivalence yields Assumption 6 for some $\gamma>0$ depending on $(1-\lambda)$ and the comparison constants, establishing item (2).
Step 2: (2) $\Rightarrow$ (1) by contradiction on the Clarke derivative. Assume item (2). Coercivity implies that, on any finite horizon, learning trajectories remain in a compact sublevel set of $\mathcal{E}$ [RockafellarWets1998, Thm. 1.9]. If some generalized signature were unbounded, we could find $(k_{n},t_{n})$, $S^{t_{n}}_{k_{n}}$ in a common compact set and $V_{n}\in\partial_{C}\Phi_{k_{n},t_{n}}(S^{t_{n}}_{k_{n}})$ with, say, $\|V_{n,x}\|\to\infty$. Choosing directions $h_{n}$ with $\|h_{n}\|=1$ and $\|V_{n,x}h_{n}\|\geq\tfrac{1}{2}\|V_{n,x}\|$, the Clarke directional derivative and chain rule along the one-step evolution force the first-order energy change to diverge:
$\limsup_{\varepsilon\downarrow 0}\frac{\mathcal{E}(S^{t_{n}+1}_{k_{n}+1}(\varepsilon))-\mathcal{E}(S^{t_{n}+1}_{k_{n}+1})}{\varepsilon}\to+\infty,$
while the dissipation inequality in Assumption 6 bounds this variation uniformly above by a linear function of $\|\delta x\|,\|\delta\theta^{0}\|,\|\delta u\|$, a contradiction (see [Clarke1990Optimization, Thms. 2.3.9 and 2.6.6]). The same argument applies to the other blocks. Hence all generalized signatures are uniformly bounded, proving item (1); the nonpositivity of generalized exponents follows from Definition 7.2. As in Option A, explicit generalized stability indices $(\tilde{\beta}_{x},\tilde{\beta}_{\theta},\tilde{\beta}_{u})$ follow from joint spectral radius / norm equivalence estimates. ∎
A.6 Proofs for section 5
Proof of Theorem 5.1.
Let the feedforward architecture satisfy the conditions in Section 5.1. Since $\sigma$ is 11–Lipschitz and differentiable a.e., the chain rule gives, for almost every $x$,
$Jf_{k}(x)=D_{k}(x)W_{k},\qquad\|D_{k}(x)\|_{2}\leq 1,$
where $D_{k}(x)$ is the diagonal (or block-diagonal) Jacobian of $\sigma$ at the pre-activation; see, e.g., standard results on a.e. differentiability of Lipschitz maps (Rademacher’s theorem) and the corresponding a.e. chain rule for compositions.
Hence, by submultiplicativity of the operator norm,
$\|Jf_{k}(x)\|_{2}\leq\|D_{k}(x)\|_{2}\,\|W_{k}\|_{2}\leq\|W_{k}\|_{2}\leq\rho<1,$
uniformly in $k$ and (a.e.) in $x$. For the depth-$L$ composition $F_{\theta}=f_{L-1}\circ\cdots\circ f_{0}$, the Jacobian at any $x$ is the ordered product
$JF_{\theta}(x)=Jf_{L-1}(X_{L-1})\cdots Jf_{0}(X_{0}),$
so again by submultiplicativity,
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|Jf_{k}(X_{k})\|_{2}\leq\rho^{L}.$
By Definition 3.1, the forward sensitivity operator coincides with the input Jacobian in this purely feedforward setting, so
$\sigma^{x}_{L,t}=\|\mathsf{S}^{x}_{L,t}\|\leq\rho^{L}\quad\text{and thus}\quad\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\leq\rho^{L}.$
Therefore, by Definition 3.1,
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)=\log\rho<0,$
where we take sequences with $T/(L+T)\to 1$ (or note that $\rho^{L}$ does not depend on $T$). Negativity of $\alpha_{x}$ yields analytic forward stability, and Theorem 4.1 then implies the existence of an analytic learning energy $\mathcal{E}$ satisfying the energy–dissipation law of Assumption 2. Since depth propagation is independent of the parameter update in the feedforward coordinate, $\mathcal{E}$ may be chosen to depend only on $(X_{L},\theta)$ (up to norm equivalence on $\mathcal{M}$). ∎
Proof of Lemma 5.2.
Fix a residual layer $X_{k+1}=X_{k}+h\,g_{k}(X_{k};\theta_{k})$ and write $G_{k}(x)=\partial_{x}g_{k}(x;\theta_{k})$. Then the Jacobian of the residual map is
$J_{k}:=\frac{\partial X_{k+1}}{\partial X_{k}}=I+hG_{k}.$
For any unit vector $v$,
$\|J_{k}v\|_{2}^{2}=\|(I+hG_{k})v\|_{2}^{2}=1+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
By the dissipativity assumption (5.2), $v^{\top}G_{k}v\leq-m$ and $\|G_{k}v\|_{2}\leq M_{g}$, hence
$\|J_{k}v\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2}.$
Taking the supremum over all unit vectors gives
$\|J_{k}\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2},$
as claimed. ∎
Proof of Theorem 5.3.
By Lemma 5.2, each residual block satisfies
$\Bigl\|\frac{\partial X_{k+1}}{\partial X_{k}}\Bigr\|_{2}\leq c_{x}(h):=\sqrt{1-2hm+h^{2}M_{g}^{2}}.$
For the depth-$L$ composition $F_{\theta}=f_{L-1}\circ\cdots\circ f_{0}$ with $f_{k}(x)=x+hg_{k}(x;\theta_{k})$, the chain rule yields
$JF_{\theta}(x)=J_{L-1}(X_{L-1})\cdots J_{0}(X_{0}),$
so by submultiplicativity,
$\|JF_{\theta}(x)\|_{2}\leq c_{x}(h)^{L}.$
Under $0<h<h_{\max}=2m/M_{g}^{2}$, we have $1-2hm+h^{2}M_{g}^{2}<1$, hence $c_{x}(h)<1$, and therefore the forward signatures satisfy
$\sigma^{x}_{k,t}\leq c_{x}(h)^{L}\qquad\text{uniformly in }k,t.$
By Definition 3.1,
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)=\log c_{x}(h)=\frac{1}{2}\log(1-2hm+h^{2}M_{g}^{2})<0,$
again taking sequences with $T/(L+T)\to 1$. Hence the representation dynamics are exponentially contractive across depth, and analytic forward stability holds. The existence of a quadratic analytic learning energy satisfying the energy–dissipation inequality of Assumption 2 then follows directly from Theorem 4.1. ∎
Proof of Corollary 5.4.
If $0<h<h_{\max}=2m/M_{g}^{2}$, then $1-2hm+h^{2}M_{g}^{2}<1$, so $c_{x}(h)<1$. By Theorem 5.3,
$\alpha_{x}=\log c_{x}(h)=\frac{1}{2}\log(1-2hm+h^{2}M_{g}^{2})<0,$
and therefore the forward sensitivity signatures are uniformly bounded and strictly contractive across depth. By Theorem 4.1, this implies analytic forward stability of the residual architecture in the sense of Definition 3.2. ∎