### v2

# Analytic and Variational Stability in Deep Learning Systems

Abstract
We propose a unified analytic and variational framework for stability in deep learning systems viewed as coupled representation–parameter dynamics. The central object is the Learning Stability Profile, which measures how infinitesimal perturbations propagate through representations, parameters, and update mechanisms along the learning trajectory.
Our main result, the Fundamental Analytic Stability Theorem, shows that uniform boundedness of these sensitivities is equivalent, up to norm equivalence, to the existence of a Lyapunov-type energy dissipating along the learning flow. In smooth regimes, this yields explicit stability exponents linking spectral norms, activation regularity, step sizes, and learning rates to contractive behavior. Classical spectral stability of feedforward networks, CFL-type conditions for residual architectures, and temporal stability laws for stochastic gradient methods follow as direct consequences.
The framework extends to non-smooth systems, including ReLU networks, proximal and projected updates, and stochastic subgradient flows, by replacing classical derivatives with Clarke generalized derivatives and smooth energies with variational Lyapunov functionals.
The resulting theory provides a unified dynamical description of stability across architectures and optimization methods, clarifying how design and training choices jointly control robustness and sensitivity to perturbations.
Keywords learning dynamics $\cdot$ stability analysis $\cdot$ Lyapunov methods $\cdot$ deep neural networks $\cdot$ stochastic gradient methods.
1 Introduction
Stability is a central concept in learning theory, but is formalized differently across numerical analysis, statistical learning, and dynamical systems. Classical notions include Lyapunov stability for discrete and continuous evolutions Hirsch et al. [2012]; Strikwerda [2004], algorithmic stability under data perturbations Bousquet and Elisseeff [2002]; Hardt et al. [2016], and Jacobian- or curvature-based criteria from control theory Sontag [1998]. These approaches address different objects and do not yield a unified analytic description of stability for modern learning systems.
We introduce a single framework that connects these notions. The central object is the Learning Stability Profile (LSP), which captures how infinitesimal perturbations propagate along the full learning flow, jointly over representations, parameters, and update dynamics. The LSP is architecture-independent and applies to smooth, nonsmooth, and stochastic learning systems, and admits continuous-time limits.
Our main result establishes an equivalence, up to norm equivalence, between boundedness of the Learning Stability Profile and the existence of a Lyapunov-type energy that dissipates along the learning trajectory. This yields a necessary and sufficient characterization of stability at the level of the learning dynamics itself. Classical stability conditions in smooth, residual, stochastic, and variational regimes are recovered as special cases Strikwerda [2004]; Kushner and Yin [2003]; Benaïm [1999]; Clarke [1990]; Rockafellar and Wets [1998].
2 Setup
We model learning as a discrete-time dynamical system on a joint state space
$Z_{k,t}=(X_{k,t},\theta_{k,t},U_{k,t}),$
with evolution
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t}).$
The maps $\Phi_{k,t}$ are assumed to be either smooth or locally Lipschitz, with generalized derivatives interpreted in the sense of Clarke when necessary, and stochastic interpretations when applicable Clarke [1990]; Kushner and Yin [2003]; Benaïm [1999].
Let $\mathcal{J}_{k,t}$ denote a (classical or generalized) Jacobian of $\Phi_{k,t}$. The sequence $(\mathcal{J}_{k,t})$ describes linearized perturbation propagation jointly across depth and training time and constitutes the Learning Stability Profile.
We quantify long-horizon sensitivity growth through analytic stability exponents. For example, the forward exponent is
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\bigl\|\partial_{Z_{0,0}}X_{k,t}\bigr\|,$
with analogous definitions for parameter and update directions. Nonpositive exponents correspond to uniform boundedness, and negative exponents to exponential decay Hirsch et al. [2012].
We also consider Lyapunov-type energies compatible with the learning flow. An analytic learning energy is a coercive, differentiable functional whose variation along trajectories satisfies a discrete dissipation inequality, in the sense of variational stability theory Rockafellar and Wets [1998].
Proposition 2.1 (Well-posedness).
The analytic stability exponents are independent of the choice of norm on the finite-dimensional state space. Uniform boundedness of linearized sensitivities is equivalent to nonpositivity of the corresponding exponents, and strict negativity implies exponential decay along sufficiently long depth–time horizons.
Proof.
All statements follow directly from equivalence of norms in finite dimensions and from the definition of the exponents as normalized logarithmic growth rates. ∎
The next section formalizes the Learning Stability Profile and establishes the equivalence between bounded analytic stability exponents and the existence of dissipative learning energies.
3 Analytic Stability Theorem
We now connect the Learning Stability Profile with Lyapunov-type energy functionals. Throughout this section we work with a learning system whose update maps are either smooth or locally Lipschitz, with generalized derivatives interpreted in the usual variational sense. We use the notation
$\mathcal{M}:=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta,$
and work up to equivalence of norms on this finite-dimensional product space.
The stability exponents summarize worst-case growth of infinitesimal perturbations in the representation, parameter, and update directions. The following theorem shows that nonpositive exponents are equivalent to the existence of a dissipative learning energy. The equivalence holds up to a change of norm and requires no convexity, stationarity, or algorithm-specific structure.
Theorem 3.1 (Fundamental Analytic Stability Theorem).
Consider a learning system whose Learning Stability Profile is well defined. Then, up to equivalence of norms on $\mathcal{M}$, the following statements are equivalent.
  1. 1.
    There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all depths $0\leq k\leq L$, all times $0\leq t\leq T$, and all horizons $L,T$,
    $\sigma^{x}_{k,t}\leq C_{x},\qquad\sigma^{\theta}_{k,t}\leq C_{\theta},\qquad\sigma^{u}_{k,t}\leq C_{u}.$
    Equivalently,
    $\alpha_{x}\leq 0,\qquad\alpha_{\theta}\leq 0,\qquad\alpha_{u}\leq 0.$
  2. 2.
    There exist a coercive, differentiable learning energy $\mathcal{E}$ and a constant $\gamma>0$ such that every trajectory of the learning flow satisfies
    $\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$
    where $R$ is linear at first order in $(\delta x,\delta\theta^{0},\delta u)$.
Moreover, when these conditions hold there exist constants $c_{1},c_{2}>0$, depending only on the regularity bounds of the learning flow, such that the stability exponents satisfy
$\alpha_{x}\leq-c_{1}\gamma,\qquad\alpha_{\theta}\leq-c_{1}\gamma,\qquad\alpha_{u}\leq-c_{2}\gamma.$
Sketch of proof.
All arguments are carried out up to equivalence of norms on $\mathcal{M}$.
Assume first that uniform bounds on the sensitivity operators hold. Then the one-step learning map is globally Lipschitz under a suitable product norm. This implies incremental stability of the learning dynamics and, by standard converse Lyapunov results for discrete-time systems, yields a coercive $C^{1}$ energy functional whose value decreases along trajectories, up to a first-order perturbation remainder. This establishes the existence of a dissipative learning energy.
Conversely, suppose such an energy exists. Linearizing the learning flow and the energy along an arbitrary trajectory yields a discrete Lyapunov inequality for the linearized system. Unbounded growth of perturbations in any coordinate direction would violate this inequality for sufficiently large horizons. A contradiction argument based on classical discrete-time Lyapunov theory implies uniform bounds on all sensitivity operators, and hence nonpositivity of the stability exponents. The quantitative bounds on the stability indices follow from norm equivalence and the dissipation rate $\gamma$. ∎
Theorem 3.1 admits natural projections isolating individual coordinates of the Learning Stability Profile and their associated exponents.
4 Residual and Feedforward Stability Laws
We now specialise the analytic stability theory to standard architectures. The goal is not to capture implementation details, but to extract simple analytic conditions that recover common stability heuristics as direct consequences of the general framework.
4.1 Feedforward networks with spectral control
Consider an $L$-layer feedforward network
$X_{k+1}=f_{k}(X_{k})=\sigma(W_{k}X_{k}+b_{k}),\qquad k=0,\dots,L-1,$
where $\sigma$ is applied elementwise and is $1$-Lipschitz. ReLU and leaky ReLU with slope in $[0,1]$ satisfy this condition.
At differentiable points, the Jacobian takes the form
$Jf_{k}(x)=D_{k}(x)W_{k},$
where $D_{k}(x)$ is diagonal with entries in $[0,1]$. Hence
$\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}.$
Let $F_{\theta}:=f_{L-1}\circ\cdots\circ f_{0}$. By the chain rule,
$JF_{\theta}(x)=Jf_{L-1}(X_{L-1})\cdots Jf_{0}(X_{0}),$
and therefore
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|W_{k}\|_{2}.$
Theorem 4.1 (Forward spectral stability for feedforward networks).
Suppose there exists $\rho\in(0,1)$ such that
$\|W_{k}\|_{2}\leq\rho\quad\text{for all }k=0,\dots,L-1.$
Then
$\sup_{x\in\mathcal{X}_{0}}\ \|JF_{\theta}(x)\|_{2}\leq\rho^{L},$
and the forward analytic exponent satisfies
$\alpha_{x}\leq\log\rho<0.$
Sketch of proof.
Each layer has Lipschitz constant at most $\rho$. Applying the chain rule yields an overall Lipschitz constant bounded by $\rho^{L}$. The exponent bound follows directly from the definition of $\alpha_{x}$. Existence of a compatible learning energy follows from Theorem 3.1. ∎
Theorem 4.1 recovers the standard principle that spectral norm control enforces stability in deep feedforward networks. Within the present framework, the forward exponent $\alpha_{x}$ plays the role of a depth-normalised logarithmic Lipschitz constant, and the condition $\rho<1$ enforces strict contraction along the representation direction.
4.2 Residual networks and CFL type stability
Residual architectures introduce an explicit step size linking depthwise propagation to time discretisations of continuous dynamical systems He et al. [2016]; Haber and Ruthotto [2017]. Within the analytic stability framework, this connection is captured quantitatively by the forward stability exponent.
Consider a residual layer
$X_{k+1}=X_{k}+h\,g_{k}(X_{k};\theta_{k}),$ (4.1)
where $h>0$ is a fixed step size and $g_{k}$ is differentiable in $x$ with Jacobian
$G_{k}(x;\theta_{k}):=\frac{\partial g_{k}}{\partial x}(x;\theta_{k}).$
Assume $g_{k}$ is globally Lipschitz and uniformly dissipative, i.e.
$\frac{G_{k}(x;\theta_{k})+G_{k}(x;\theta_{k})^{\top}}{2}\preceq-mI,\qquad\|G_{k}(x;\theta_{k})\|_{2}\leq M_{g},$ (4.2)
for all admissible $(x,\theta_{k})$ and all $k$. Such conditions are standard in the stability analysis of explicit discretisations of dissipative systems Strikwerda [2004].
At each layer,
$\frac{\partial X_{k+1}}{\partial X_{k}}=I+h\,G_{k}(X_{k};\theta_{k}).$
Lemma 4.2 (One step residual Lipschitz factor).
Under assumption (4.2), every residual layer satisfies
$\Bigl\|\frac{\partial X_{k+1}}{\partial X_{k}}\Bigr\|_{2}^{2}\leq 1-2hm+h^{2}M_{g}^{2}$
for all admissible $(X_{k},\theta_{k})$.
Proof.
For any unit vector $v$,
$\|(I+hG_{k})v\|_{2}^{2}=1+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|_{2}^{2}.$
The dissipativity and Lipschitz bounds imply $v^{\top}G_{k}v\leq-m$ and $\|G_{k}v\|_{2}^{2}\leq M_{g}^{2}$, yielding the claim. ∎
Define the one-step contraction factor
$c_{x}(h):=\sqrt{1-2hm+h^{2}M_{g}^{2}}.$
Theorem 4.3 (Analytic stability of residual networks).
Consider the residual architecture (4.1) under assumption (4.2). If the step size satisfies
$0<h<h_{\max},\qquad h_{\max}:=\frac{2m}{M_{g}^{2}},$ (4.3)
then $c_{x}(h)<1$ and the forward analytic exponent satisfies
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0.$
Sketch of proof.
Lemma 4.2 gives a uniform bound $c_{x}(h)$ on each layer. Composition over depth yields a forward Jacobian norm bounded by $c_{x}(h)^{L}$. The exponent bound follows directly from the definition of the analytic exponent given in Section 2.
Condition (4.3) is a discrete stability restriction identical in structure to a Courant-Friedrichs-Lewy condition for explicit time stepping of dissipative systems Strikwerda [2004]. The analytic stability framework therefore formalises the heuristic that residual blocks must be either sufficiently small in step size or sufficiently dissipative to ensure stable depthwise propagation.
Corollary 4.4 (CFL-type stability condition for residual networks).
Under the hypotheses of Theorem 4.3, the residual architecture is forward analytically stable whenever
$0<h<h_{\max},\qquad h_{\max}=\frac{2m}{M_{g}^{2}}.$
In this regime,
$\alpha_{x}\leq\frac{1}{2}\log\bigl(1-2hm+h^{2}M_{g}^{2}\bigr)<0.$
5 Temporal stability of stochastic gradient methods
We consider stochastic gradient updates
$\theta^{t+1}=\theta^{t}-\eta_{t}\,G(\theta^{t},u^{t}),$ (5.1)
under standard strong convexity, smoothness, and bounded noise assumptions Nesterov [2004]; Kushner and Yin [2003].
We assume that the stochastic gradient map is uniformly Lipschitz in both parameters and update randomness. Namely, there exist constants $L_{G},L_{u}^{G}\geq 0$ such that
$\|G(\theta_{1},u)-G(\theta_{2},u)\|\leq L_{G}\|\theta_{1}-\theta_{2}\|,\qquad\|G(\theta,u_{1})-G(\theta,u_{2})\|\leq L_{u}^{G}\|u_{1}-u_{2}\|.$ (5.2)
Assumption 1 (Stochastic gradient regime).
The objective $\mathcal{L}$ is $\mu$–strongly convex and LL–smooth. The stochastic gradient satisfies
$\mathbb{E}[G(\theta,u)\mid\theta]=\nabla\mathcal{L}(\theta),\qquad\mathbb{E}\!\left[\|G(\theta,u)-\nabla\mathcal{L}(\theta)\|^{2}\mid\theta\right]\leq\sigma_{0}^{2}+\sigma_{1}^{2}\|\theta-\theta^{\star}\|^{2}.$
Moreover, $G$ is Lipschitz in $\theta$ and $u$ as in (5.2).
5.1 Mean square Lyapunov recursion
Let $e^{t}:=\theta^{t}-\theta^{\star}$. Standard expansion yields the conditional recursion
$\mathbb{E}\bigl[\|e^{t+1}\|^{2}\mid\theta^{t}\bigr]\leq\bigl(1-2\eta_{t}\mu+\eta_{t}^{2}(L^{2}+\sigma_{1}^{2})\bigr)\|e^{t}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2}.$ (5.3)
Lemma 5.1 (Mean square Lyapunov recursion for SGD).
Under Assumption 1, the SGD update (5.1) satisfies
$\mathbb{E}\bigl[\|\theta^{t+1}-\theta^{\star}\|^{2}\mid\theta^{t}\bigr]\leq\|\theta^{t}-\theta^{\star}\|^{2}-\eta_{t}\bigl(2\mu-\eta_{t}(L^{2}+\sigma_{1}^{2})\bigr)\|\theta^{t}-\theta^{\star}\|^{2}+\eta_{t}^{2}\sigma_{0}^{2}.$
5.2 Temporal analytic stability for constant and decreasing step sizes
We now translate the Lyapunov recursion (5.3) into bounds on the parametric and temporal analytic exponents of the Learning Stability Profile.
Theorem 5.2 (Analytic temporal stability of SGD with constant step size).
Let Assumption 1 hold and suppose $\eta_{t}\equiv\eta$ is constant. Assume
$0<\eta<\eta_{\max},\qquad\eta_{\max}:=\frac{2\mu}{L^{2}+\sigma_{1}^{2}}.$ (5.4)
Then $q(\eta):=1-2\eta\mu+\eta^{2}(L^{2}+\sigma_{1}^{2})$ satisfies $0<q(\eta)<1$ and the following hold.
  1. 1.
    The quadratic energy $\mathcal{E}(\theta)=\|\theta-\theta^{\star}\|^{2}$ satisfies
    $\mathbb{E}\bigl[\mathcal{E}(\theta^{t+1})\bigr]\leq q(\eta)\,\mathbb{E}\bigl[\mathcal{E}(\theta^{t})\bigr]+\eta^{2}\sigma_{0}^{2}$
    for all $t$.
  2. 2.
    The parametric analytic exponent satisfies
    $\alpha_{\theta}\leq\log q(\eta)<0.$
  3. 3.
    The temporal analytic exponent satisfies $\alpha_{u}\leq 0$. More precisely, there exists $C_{u}>0$, depending only on $(L_{G},L_{u}^{G},\mu,L)$ and on the chosen product norm on $\mathcal{M}$, such that
    $\|\mathsf{S}^{u}_{t}\|\leq C_{u}\,\eta\,L_{u}^{G}\qquad\text{for all }t.$
Sketch of proof.
Part (1) follows directly from (5.3). The linear contraction of the mean update map implies exponential forgetting of initial conditions in mean square, yielding part (2). Part (3) follows from differentiating the update with respect to $u$ and applying standard input–output bounds for linear time-varying systems [Khalil, 2002, Chapter 4]. ∎
With constant step size, stability is not asymptotic in the deterministic sense due to persistent noise. The exponent $\alpha_{\theta}$ nevertheless captures exponential decay of initialization effects, while $\alpha_{u}$ quantifies uniform sensitivity to stochastic perturbations.
Theorem 5.3 (Analytic temporal stability with decreasing step size).
Let Assumption 1 hold and suppose that the step sizes satisfy
$\eta_{t}>0,\qquad\sum_{t=0}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=0}^{\infty}\eta_{t}^{2}<\infty.$
Then the following statements hold.
  1. 1.
    The iterates $\theta^{t}$ converge almost surely to $\theta^{\star}$, and the energy $\mathcal{E}(\theta^{t})$ converges almost surely and is nonincreasing in expectation.
  2. 2.
    The parametric analytic exponent satisfies
    $\alpha_{\theta}\leq 0.$
  3. 3.
    The temporal analytic exponent satisfies
    $\alpha_{u}\leq 0,$
    and the temporal stability signatures remain uniformly bounded.
Sketch of proof.
The argument follows the classical ODE method for stochastic approximation Kushner and Yin [2003]; Benaïm [1999]. Almost sure convergence and boundedness of sensitivities imply nonexpansiveness of the linearised dynamics, yielding $\alpha_{\theta}\leq 0$ and $\alpha_{u}\leq 0$. ∎
Theorems 5.2 and 5.3 embed stochastic gradient methods within the analytic stability framework. The learning rate determines whether the parametric and temporal exponents are strictly negative or merely nonpositive, while the Learning Stability Profile provides a unified description of robustness to initialization and stochastic update perturbations.
6 Variational Extension to Non-smooth Learning Systems
Many learning architectures involve non-smooth components, including ReLU-type activations, nonsmooth losses, clipping, quantisation, and proximal or projected updates. This section extends the analytic stability framework to such systems.
The extension replaces classical Jacobians with Clarke generalized derivatives and smooth Lyapunov energies with variational energies. The central equivalence established earlier remains valid: bounded stability signatures are equivalent to the existence of a dissipative learning energy, with derivatives interpreted in the Clarke sense Clarke [1990]; Clarke et al. [1998]; Rockafellar and Wets [1998].
6.1 Locally Lipschitz learning flows
We replace differentiability assumptions by local Lipschitz continuity.
Assumption 2 (Locally Lipschitz learning flow).
For each index $(k,t)$, the map
$\Phi_{k,t}\colon(x,\theta,u)\mapsto\bigl(f_{k}(x;\theta_{k}),\mathcal{A}_{t}(\theta,u)\bigr)$
is locally Lipschitz. There exist constants
$L^{\mathrm{C}}_{\mathrm{fwd}},\qquad L^{\mathrm{C}}_{\mathrm{par}},\qquad L^{\mathrm{C}}_{\mathrm{upd}}$
such that for every $(x,\theta,u)$ and every Clarke generalized Jacobian
$V\in\partial_{C}\Phi_{k,t}(x,\theta,u),$
the corresponding blocks satisfy
$\|V_{x}\|\leq L^{\mathrm{C}}_{\mathrm{fwd}},\qquad\|V_{\theta}\|\leq L^{\mathrm{C}}_{\mathrm{par}},\qquad\|V_{u}\|\leq L^{\mathrm{C}}_{\mathrm{upd}}.$
Local Lipschitz continuity guarantees that the Clarke generalized Jacobian is a nonempty compact convex set and varies upper semicontinuously Clarke [1990]. This assumption covers common non-smooth components used in learning systems, including piecewise affine activations, proximal operators, and convex projections Rockafellar and Wets [1998]; Bauschke and Combettes [2017].
Under Assumption 2, the learning dynamics are described by the difference inclusion
$S^{t+1}_{k+1}\in\Phi_{k,t}(S^{t}_{k}),\qquad S^{t}_{k}=(X^{t}_{k},\theta^{t}),$
with existence of trajectories ensured by standard results for discrete-time locally Lipschitz systems Clarke [1990].
6.2 Generalized Learning Stability Profile
We extend the Learning Stability Profile by measuring sensitivities using Clarke generalized Jacobians.
Definition 6.1 (Generalized Learning Stability Profile).
Under Assumption 2, the generalized stability signature at $(k,t)$ and state $S^{t}_{k}$ is
$\mathrm{gLSP}(k,t;S^{t}_{k})=\bigl(\sigma^{x,C}_{k,t},\sigma^{\theta,C}_{k,t},\sigma^{u,C}_{k,t}\bigr),$
where
$\sigma^{x,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{x}\|,\quad\sigma^{\theta,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{\theta}\|,\quad\sigma^{u,C}_{k,t}:=\sup_{V\in\partial_{C}\Phi_{k,t}(S^{t}_{k})}\|V_{u}\|.$
When $\Phi_{k,t}$ is differentiable at $S^{t}_{k}$, the generalized profile coincides with the smooth Learning Stability Profile. In general, it captures worst-case local amplification over all active linearizations of the non-smooth learning flow.
6.3 Generalized analytic exponents
Definition 6.2 (Generalized analytic exponents).
Let $L$ and $T$ denote network depth and training horizon. The generalized analytic exponents are defined by
$\tilde{\alpha}_{x}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{x,C}_{k,t},$
$\tilde{\alpha}_{\theta}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{\theta,C}_{k,t},$
$\tilde{\alpha}_{u}$ $:=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\sigma^{u,C}_{k,t}.$
Definition 6.3 (Generalized analytic stability).
The learning system is generalized analytically stable if
$\tilde{\alpha}_{x}\leq 0,\qquad\tilde{\alpha}_{\theta}\leq 0,\qquad\tilde{\alpha}_{u}\leq 0.$
Negative generalized exponents correspond to contractive generalized linearized dynamics, while nonpositive values correspond to uniform boundedness of sensitivities.
6.4 Variational energies and dissipation
Assumption 3 (Variational learning energy).
A variational learning energy is a proper lower semicontinuous functional
$\mathcal{E}:\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}\cup\{+\infty\},$
finite on all admissible states and satisfying coercivity bounds
$c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr),$
for constants $0<c_{1}\leq c_{2}<\infty$.
Assumption 4 (Variational energy dissipation).
There exist a variational learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$, linear at first order in $(\delta x,\delta\theta^{0},\delta u)$, such that for every training step $t$,
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u).$
6.5 Fundamental Variational Stability Theorem
Theorem 6.4 (Fundamental Variational Stability Theorem).
Let the flow maps $\Phi_{k,t}$ satisfy Assumption 2. Up to equivalence of norms on $\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta$, the following are equivalent.
  1. 1.
    There exist constants $C_{x},C_{\theta},C_{u}\geq 1$ such that for all $k,t$,
    $\sigma^{x,C}_{k,t}\leq C_{x},\qquad\sigma^{\theta,C}_{k,t}\leq C_{\theta},\qquad\sigma^{u,C}_{k,t}\leq C_{u}.$
    Equivalently,
    $\tilde{\alpha}_{x}\leq 0,\qquad\tilde{\alpha}_{\theta}\leq 0,\qquad\tilde{\alpha}_{u}\leq 0.$
  2. 2.
    There exists a variational learning energy $\mathcal{E}$ satisfying Assumptions 3 and 4.
Sketch of proof.
Uniform bounds on generalized stability signatures imply incremental stability of the associated difference inclusion under a suitable product norm. Converse Lyapunov results for nonsmooth discrete-time systems yield a coercive variational energy satisfying the dissipation inequality Clarke [1990]; Clarke et al. [1998].
Conversely, coercivity and dissipation preclude unbounded growth of generalized perturbations. Selecting elements of the Clarke generalized Jacobian and applying generalized chain rules yields uniform bounds on generalized sensitivities, implying nonpositivity of the generalized analytic exponents. ∎
The theorem shows that analytic stability is preserved when passing from smooth to non-smooth learning systems. Generalized analytic exponents retain the same interpretation as in the smooth case, quantifying contractivity or boundedness of learning dynamics with respect to inputs, parameters, and update mechanisms.
6.6 Examples: piecewise linear networks and proximal flows
We illustrate how the variational framework applies to two representative non-smooth systems by verifying uniform bounds on generalized stability signatures and the existence of a dissipative variational energy. Detailed derivations follow the same pattern as in Sections 4, 5, and 6 and are therefore omitted.
Piecewise linear networks.
Consider a feedforward architecture
$X_{k+1}=f_{k}(X_{k}):=\sigma(W_{k}X_{k}+b_{k}),\qquad k=0,\dots,L-1,$
where $\sigma$ is piecewise linear with slopes in $[0,1]$. Each $f_{k}$ is globally Lipschitz and differentiable almost everywhere. At differentiable points
$Jf_{k}(x)=D_{k}(x)W_{k},$
with $D_{k}(x)$ diagonal and entries in $[0,1]$. The Clarke generalized Jacobian satisfies
$\partial_{C}f_{k}(x)=\{DW_{k}\mid D\text{ diagonal},\ D_{ii}\in[0,1]\},$
so for any compatible operator norm
$\sup_{V\in\partial_{C}f_{k}(x)}\|V\|\leq\|W_{k}\|.$
If $\|W_{k}\|_{2}\leq\rho<1$ for all $k$, then Assumption 2 holds with $L^{\mathrm{C}}_{\mathrm{fwd}}\leq\rho$ and finite $(L^{\mathrm{C}}_{\mathrm{par}},L^{\mathrm{C}}_{\mathrm{upd}})$. In particular,
$\sigma^{x,C}_{k,t}\leq\rho\quad\text{for all }(k,t),\qquad\tilde{\alpha}_{x}\leq\log\rho<0.$
For fixed parameters, the quadratic functional
$\mathcal{E}_{\mathrm{fwd}}(X_{0},\dots,X_{L},\theta):=\tfrac{1}{2}\|X_{L}\|^{2}$
is proper, lower semicontinuous, and coercive in $X_{L}$. Layerwise Lipschitz bounds yield a discrete Lyapunov inequality of the form Assumption 4 for a modified quadratic energy, with rate controlled by $1-\rho^{2}$. Hence the hypotheses of Theorem 6.4 are satisfied, and the generalized forward exponent can be read equivalently from Jacobian bounds or energy decay.
Proximal and projected flows.
Consider $\Phi(\theta)=\mathcal{L}(\theta)+R(\theta)$, where $\mathcal{L}$ has $L$-Lipschitz gradient and $R$ is proper, lower semicontinuous, and convex. The proximal gradient update
$\theta^{t+1}=\operatorname{prox}_{\eta R}\bigl(\theta^{t}-\eta\nabla\mathcal{L}(\theta^{t})\bigr)$
can be written as $T_{\eta}\circ S_{\eta}$, where $S_{\eta}(\theta)=\theta-\eta\nabla\mathcal{L}(\theta)$ and $T_{\eta}=\operatorname{prox}_{\eta R}$. For $0<\eta<2/L$, $S_{\eta}$ is nonexpansive and $T_{\eta}$ is firmly nonexpansive, hence $T_{\eta}\circ S_{\eta}$ is globally Lipschitz with constant at most one. Assumption 2 therefore holds with
$\sigma^{\theta,C}_{t}\leq 1\quad\text{for all }t,\qquad\tilde{\alpha}_{\theta}\leq 0.$
On the energy side, $\Phi$ itself is a variational learning energy. Under standard regularity conditions, the proximal gradient step satisfies
$\Phi(\theta^{t+1})-\Phi(\theta^{t})\leq-c_{\eta}\|\theta^{t+1}-\theta^{t}\|^{2},$
for some $c_{\eta}>0$, which is a special case of Assumption 4. The Fundamental Variational Stability Theorem 6.4 therefore applies to proximal and projected gradient flows, including the case where $R$ is the indicator of a closed convex set.
These examples illustrate that architectural and optimisation-induced non-smoothness fit naturally within the same stability grammar.
7 Unified Interpretation
The results of this paper define an explicit stability grammar for learning systems. The (generalized) Learning Stability Profile characterises how perturbations in inputs, parameters, and update mechanisms propagate through the coupled representation–parameter flow. The Fundamental Analytic Stability Theorem and its variational counterpart show that uniform control of these sensitivities is equivalent to the existence of an energy functional that dissipates along learning trajectories, up to equivalence of norms on the joint state space.
7.1 Stability as a joint property of architecture and optimisation
In previous sections, learning is modelled as a discrete-time dynamical system on a joint state space containing all layer representations and parameters. The Learning Stability Profile records operator norms of directional sensitivities derived from the Jacobian or generalized Jacobian of the coupled flow. The analytic and generalized exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ and $(\tilde{\alpha}_{x},\tilde{\alpha}_{\theta},\tilde{\alpha}_{u})$ summarise the asymptotic growth or decay of these sensitivities along depth and training time.
The Fundamental Analytic Stability Theorem 3.1 and the Fundamental Variational Stability Theorem 6.4 show that, in both smooth and non-smooth regimes, nonpositive exponents are equivalent to the existence of an energy functional satisfying a discrete dissipation inequality. The choice between smooth and variational formulations is dictated by regularity: $C^{1}$ systems admit a Jacobian-based analysis, while locally Lipschitz systems require Clarke generalized derivatives and lower semicontinuous energies. On their overlap, the two formulations coincide, since the Clarke generalized Jacobian reduces to the classical Jacobian on a dense set and the associated exponents agree up to norm equivalence Clarke [1990]; Rockafellar and Wets [1998].
This joint perspective links local analytic quantities, such as spectral norms, Lipschitz constants, and step sizes, to global dynamical behaviour through stability exponents and energies. Architecture and optimisation are treated symmetrically as components of a single coupled flow. The feedforward, residual, stochastic gradient, and variational stability laws derived in Sections 4, 5, and 6 are specialisations of this unified criterion.
7.2 Connection with classical stability notions
The analytic–variational framework unifies stability concepts across several communities. In numerical analysis and control, Lyapunov theory relates contractivity and negative exponents to strict Lyapunov functions Khalil [2002], while CFL-type conditions ensure stability of explicit schemes for dissipative systems Strikwerda [2004]. In optimisation, firm nonexpansiveness and averagedness yield contractive behaviour for proximal and splitting methods Bauschke and Combettes [2017]; Bolte et al. [2014]; Davis and Drusvyatskiy [2019]. In statistical learning, algorithmic stability bounds generalisation error via insensitivity to perturbations Bousquet and Elisseeff [2002]; Hardt et al. [2016]; Bottou et al. [2018].
Within the present framework, the Learning Stability Profile acts as a multi-directional Lyapunov exponent for the coupled learning flow. The analytic exponents reduce to spectral bounds for linear systems and to classical Lyapunov exponents in continuous-time limits, while the associated energy functionals generalise discrete Lyapunov functions to joint representation–parameter dynamics. Proximal and projected methods correspond to cases where the variational energy coincides with the objective, while algorithmic stability corresponds to nonpositive parametric and temporal exponents.
The feedforward and residual stability laws recover discrete counterparts of stability results for neural ordinary differential equations Haber and Ruthotto [2017]; Ruthotto and Haber [2020]. The CFL-type step size conditions correspond to admissible time steps for explicit discretisations of dissipative flows, while the stochastic gradient stability law recovers classical stochastic approximation conditions Kushner and Yin [2003]; Benaïm [1999] in exponent form.
7.3 Design rules and trade offs
The explicit dependence of stability exponents on spectral norms, Lipschitz constants, dissipation rates, and learning rates yields concrete design rules. For feedforward networks, the forward exponent is controlled by layerwise spectral bounds, while for residual architectures it is governed by the product of the residual step size and the Jacobian bound of each block. The CFL-type constraint in Corollary 4.4 therefore acts as a design condition linking architectural scale and step size.
For stochastic gradient methods, the temporal stability law expresses the parametric exponent in terms of learning rate, curvature, and gradient noise. The effective contraction factor $q(\eta)$ in Theorem 5.2 refines deterministic stability criteria by incorporating stochastic effects, explaining the sensitivity of training to learning-rate schedules in high-variance regimes Bottou et al. [2018].
The framework also clarifies accuracy–stability trade offs. Impossibility results for certain expressive architectures Colbrook et al. [2021] correspond to regimes where the forward exponent must become positive as accuracy is pushed beyond a threshold. The analytic exponents thus provide a principled way to express stability constraints in terms of geometry and spectral scale rather than empirical performance alone.
7.4 Robustness, generalisation, and explanations
The analytic and generalized exponents connect stability to robustness and generalisation. Algorithmic stability theory links small generalisation gaps to insensitivity of learning procedures Bousquet and Elisseeff [2002]; Hardt et al. [2016]. In the present framework, bounded parametric and temporal exponents imply that perturbations in training data or stochasticity induce uniformly bounded changes in parameters and representations, as quantified by the Learning Stability Profile.
The same mechanism governs the stability of explanation methods. Explanation functionals based on gradients or saliency inherit their sensitivity from network Jacobians and from the stability of the training procedure Fel et al. [2022]. When forward or parametric exponents are close to zero or positive, small perturbations may be amplified, leading to unstable explanations even when prediction error is low. Negative exponents instead imply quantitative robustness of both predictions and a broad class of gradient-based explanations. While interpretability also involves fidelity and domain alignment Abusitta et al. [2024], the analytic–variational framework isolates stability as a structural property governed by architecture and optimisation.
7.5 Scope and limitations
The framework applies to finite-dimensional state spaces and discrete-time learning dynamics. The analytic results require global bounds on classical or generalized Jacobians, while the variational results require lower semicontinuous energies with quadratic coercivity along trajectories of interest. These assumptions cover many standard architectures and optimisers, including spectrally controlled feedforward networks, residual networks with bounded Jacobians, stochastic gradient methods under classical conditions, and proximal or projected flows.
They do not cover all regimes encountered in practice. Strongly adaptive optimisers, non-stationary data, heavy-tailed noise, or unbounded parameter growth may violate global bounds or admit only local or state-dependent energies Bottou et al. [2018]. The theory is therefore structural rather than fully quantitative: the derived inequalities relate exponents, regularity constants, and dissipation rates, but are not claimed to be sharp in high-dimensional settings.
The analytic and generalized exponents should thus be viewed as organising quantities capturing stability at the level of mechanisms and design rules, rather than as precise numerical predictors for specific large models. They unify Lyapunov stability, CFL-type conditions, proximal contraction, and algorithmic stability within a single analytic–variational framework.
8 Conclusion
We introduced an analytic and variational framework for stability in deep learning systems based on the Learning Stability Profile, which quantifies how perturbations propagate through coupled representation and parameter dynamics. The Fundamental Analytic Stability Theorem and its variational extension show that bounded sensitivities are equivalent to the existence of dissipative Lyapunov-type energies, without requiring convexity or algorithm-specific structure.
Within this framework, classical stability conditions for feedforward, residual, and stochastic gradient methods arise as special cases, while the theory extends naturally to non-smooth systems through Clarke generalized derivatives and variational energies. Stability is therefore expressed through a small set of exponents directly linked to architectural and optimization choices.
Future directions include continuous-time limits, geometric formulations of learning dynamics, and practical estimation of stability indices in large-scale models. Overall, the framework provides a structural basis for principled architecture and optimization design grounded in analytic stability guarantees.
Appendix A Technical Proofs
This appendix contains proofs of results stated in the main text. Standard facts repeatedly used include equivalence of norms in finite dimensions, operator norm bounds, and classical Lyapunov arguments; see [Horn and Johnson, 2013, Section 5.4] and [Khalil, 2002, Chapter 4].
Only arguments specific to the present framework are detailed. Classical steps are cited rather than reproved.
A.1 Proof of Theorem 3.1
Proof.
Let
$\mathcal{M}=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta$
be equipped with an arbitrary norm. All norms on $\mathcal{M}$ are equivalent.
Denote by $\Psi_{t}$ the one-step learning map $S^{t}_{\cdot}\mapsto S^{t+1}_{\cdot}$ obtained by composing depthwise representation updates with parameter updates at time $t$.
[…]
Uniform bounds on the signatures $\sigma^{x}_{k,t}$, $\sigma^{\theta}_{k,t}$, and $\sigma^{u}_{k,t}$ imply uniform bounds on derivatives of $\Psi_{t}$ after passing to an equivalent product norm. Hence each $\Psi_{t}$ is globally Lipschitz.
Nonpositive analytic exponents mean perturbations do not grow uniformly over depth and time, so the coupled dynamics are incrementally stable. Converse Lyapunov theorems for incrementally stable discrete systems then yield a $C^{1}$ incremental Lyapunov function Angeli [2002]; Khalil [2002] satisfying quadratic bounds and strict decrease along trajectories.
Fixing a reference trajectory converts this incremental Lyapunov function into a coercive learning energy $\mathcal{E}$ whose one-step variation satisfies the dissipation inequality of item (2). First-order perturbation terms produce the remainder $R$.
[…]
Assume a coercive differentiable energy satisfies the dissipation inequality. Setting perturbations to zero yields strict decay along trajectories. Coercivity bounds state norms of relevant coordinates are controlled by $\mathcal{E}$, so growth of perturbations would contradict energy decay.
Linearizing the dynamics yields a linear time-varying system whose transition operators must therefore remain uniformly bounded [Khalil, 2002, Chapter 4]. Projection onto coordinate blocks gives uniform bounds on the sensitivity operators, hence $\alpha_{x},\alpha_{\theta},\alpha_{u}\leq 0$.
Quantitative index bounds follow from quadratic comparison constants and norm equivalence. ∎
A.2 Proofs for Section 4
Proof of Theorem 4.1.
Proof.
Each layer satisfies $\|Jf_{k}(x)\|_{2}\leq\|W_{k}\|_{2}\leq\rho$ since $\sigma$ is $1$-Lipschitz. Chain rule and submultiplicativity yield
$\|JF_{\theta}(x)\|_{2}\leq\prod_{k=0}^{L-1}\|W_{k}\|_{2}\leq\rho^{L}.$
Taking logarithmic growth rates gives $\alpha_{x}\leq\log\rho<0$. ∎
Proof of Lemma 4.2.
Proof.
For unit $v$,
$\|(I+hG_{k})v\|^{2}=1+2h\,v^{\top}G_{k}v+h^{2}\|G_{k}v\|^{2}.$
Dissipativity gives $v^{\top}G_{k}v\leq-m$, and Lipschitz bounds yield $\|G_{k}v\|^{2}\leq M_{g}^{2}$, proving the claim. ∎
Proof of Theorem 4.3.
Proof.
Lemma 4.2 yields contraction factor $c_{x}(h)=\sqrt{1-2hm+h^{2}M_{g}^{2}}$ per layer. Composing over depth gives contraction $c_{x}(h)^{L}$. Condition $h<h_{\max}$ ensures $c_{x}(h)<1$, yielding $\alpha_{x}<0$. ∎
A.3 Proofs for Section 5
Proof of Lemma 5.1.
Proof.
Let $e^{t}=\theta^{t}-\theta^{\star}$. Expanding $\|e^{t+1}\|^{2}$ using update $\theta^{t+1}=\theta^{t}-\eta_{t}G(\theta^{t},u^{t})$ gives
$\|e^{t+1}\|^{2}=\|e^{t}\|^{2}-2\eta_{t}e^{t\top}G(\theta^{t},u^{t})+\eta_{t}^{2}\|G(\theta^{t},u^{t})\|^{2}.$
Taking conditional expectations and using unbiasedness, strong convexity, and variance bounds yields the recursion. ∎
Proof of Theorem 5.2.
Proof.
The Lyapunov recursion implies
$\mathbb{E}\|e^{t+1}\|^{2}\leq q(\eta)\mathbb{E}\|e^{t}\|^{2}+\eta^{2}\sigma_{0}^{2}.$
Contraction of coupled trajectories yields exponential decay of initialization sensitivity, so $\alpha_{\theta}<0$. Input–output bounds for the linearized recursion give bounded temporal sensitivity, hence $\alpha_{u}\leq 0$. ∎
Proof of Theorem 5.3.
Proof.
Standard stochastic approximation arguments using Robbins–Siegmund theorem yield almost sure convergence Kushner and Yin [2003]; Benaïm [1999]. Coupled recursion bounds imply bounded sensitivity operators, so $\alpha_{\theta},\alpha_{u}\leq 0$. ∎
A.4 Proofs for Section 6
Lemma A.1 (Smooth reduction).
If $\Phi_{k,t}$ is differentiable at a state, the Clarke generalized Jacobian reduces to the classical Jacobian, so generalized and analytic profiles coincide.
Proof.
This is standard Clarke theory; see [Clarke, 1990, Proposition 2.6.2]. ∎
Proof of Theorem 6.4.
Proof.
Uniform bounds on generalized signatures imply incremental stability of the difference inclusion. Converse Lyapunov theorems for nonsmooth systems yield a proper lower semicontinuous Lyapunov function satisfying dissipation inequalities Clarke et al. [1998].
Conversely, coercivity and dissipation prevent unbounded growth of generalized linearizations. Clarke chain rules then imply uniform bounds on generalized Jacobians, giving nonpositive generalized exponents.
Strict dissipation yields strict negativity of indices through standard spectral-radius arguments. ∎